### v1

# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

Abstract
The convergence of two technological trajectories, large language models that automate cognitive labor and deep reinforcement learning agents that automate physical labor, implies the near-complete elimination of human employment as historically understood. The universal approximation theorem establishes that any continuous function can be represented by a sufficiently expressive neural network; foundational results in deep reinforcement learning establish that optimal policies exist for sequential decision tasks. Together, these results imply that all labor, white-collar and blue-collar, is in principle automatable, though significant engineering challenges remain. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care rather than corporate extraction. Synthesizing anarchist political theory, commons-based peer production, and critical technology studies, we propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons built on existing open-source technologies. We demonstrate that anarcho-accelerationism functions as a bridge-building framework compatible with liberalism, socialism, environmentalism, feminism, cooperativism, and the hacker ethic. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net shows that commons-based systems already operate at scale. We further argue that full automation renders money obsolete, making Universal Basic Income an insufficient response; we propose instead Universal Desired Resources (UDR), a post-monetary design principle where every person requests what they need from the robotic commons, constrained only by ecological sustainability. Drawing on prior work demonstrating the independence of phenomenal consciousness from computational intelligence, we establish that delegating all labor to non-conscious machines is not exploitation but care at civilizational scale, and that moral policy for the commons can be studied and refined through deep reinforcement learning. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
Keywords: commons-based peer production; federated systems; cooperative platforms; sociotechnical design; mutual aid; digital sovereignty; open-source AI; mesh networks; platform cooperativism; democratic technology; universal approximation; deep reinforcement learning; full automation; universal desired resources; machine consciousness; computational ethics
1 Introduction: The Convergence Window
We live in a paradox. Humanity possesses the technical capacity to feed, house, heal, educate, and connect every person on Earth (Bookchin, 1971). Solar energy is now the cheapest electricity source in history (IRENA, 2023). A single modern smartphone carries more computational power than the Apollo mission control. Open-source software runs 96% of the world’s top web servers (Linux Foundation, 2023). And yet 735 million people face chronic hunger (FAO et al., 2023), three individuals hold more wealth than the bottom half of the United States (Oxfam International, 2023), and the six largest technology corporations exercise more influence over global communication than any state in history (Zuboff, 2019).
This is not a resource problem. It is an architecture problem.
The dominant architectures of our sociotechnical systems (centralized cloud platforms, extractive business models, proprietary codebases, fossil-fuel energy grids, top-down governance) were designed to concentrate power. They are spectacularly effective at this task. But they are not the only architectures available to us. A growing body of evidence from computer science, institutional economics, cooperative studies, and political ecology suggests that decentralized, commons-based, federated architectures can achieve equal or superior performance in coordination, resilience, innovation, and human well-being (Benkler, 2006; Ostrom, 1990; Scholz, 2016).
This paper introduces anarcho-accelerationism: a framework for sociotechnical system design that takes this evidence seriously. The name is deliberately provocative, but the substance is rigorous. The prefix anarcho- refers not to chaos but to the 150-year tradition of anarchist political philosophy, from Kropotkin’s empirical study of mutual aid in nature (Kropotkin, 1902) to Bookchin’s ecological municipalism (Bookchin, 2005) to Ostrom’s Nobel Prize-winning work on commons governance (Ostrom, 1990). The core principle is self-organization without domination. The suffix -accelerationism refers not to reckless technological enthusiasm but to the critical insight that current productive forces, if redirected from accumulation toward liberation, are sufficient to abolish artificial scarcity and enable universal flourishing (Srnicek and Williams, 2015). The core principle is technological sufficiency for post-scarcity.
The synthesis yields a design philosophy centered on three values that may sound naive to computer scientists trained in optimization and efficiency, but which we will show are both technically implementable and empirically validated. First, care: systems should be designed to sustain life, relationships, and communities rather than to maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017). Second, love: not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956). Third, the common good: resources that are essential for human dignity (energy, food, knowledge, communication, code) should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007).
This urgency is not abstract. As we demonstrate in Section˜2, foundational theorems in machine learning (the universal approximation theorem and the convergence results of deep reinforcement learning) guarantee that all human labor, cognitive and physical, is in principle automatable. The empirical evidence, from LLM-powered coding agents to humanoid robots, confirms that this automation is not a distant prospect but an ongoing process. The question is no longer whether jobs will disappear, but who will control the machines that replace them.
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜3), while offering a technical architecture (Section˜4) grounded in empirical evidence (Section˜5). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
2 The Automation Imperative: Why This Framework, Why Now
The urgency of anarcho-accelerationism stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. We ground this claim in two foundational results from machine learning theory and their observable consequences.
The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. The implication is significant: every task that can be described as a continuous mapping from inputs to outputs can in principle be represented by a neural network. We note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. The theorem applies only to continuous functions on compact subsets, and the number of training samples required (sample complexity) may be prohibitively large. Our claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
The convergence theorems for deep reinforcement learning (DRL) (Sutton and Barto, 2018; Mnih et al., 2015) establish that an agent interacting with any environment describable as a Markov Decision Process (MDP) can learn an optimal policy through observation of states, actions, and rewards. For any MDP $(S,A,P,R,\gamma)$ with finite state and action spaces, Q-learning converges to the optimal action-value function $Q^{*}$ given sufficient exploration (Watkins and Dayan, 1992). With deep neural networks as function approximators (DQN and its successors), Q-learning has achieved remarkable empirical success in continuous, high-dimensional state spaces (Mnih et al., 2015), though we note that formal convergence guarantees do not extend from the tabular case to function approximation in general: the combination of bootstrapping, off-policy learning, and function approximation (the “deadly triad”) can cause divergence (Sutton and Barto, 2018). The practical successes summarized in Table˜1 are therefore empirical achievements, not deductive consequences of the tabular convergence theorem. Combined with imitation learning, where policies are acquired by observing human demonstrations rather than through explicit reward signals (Schaal, 1999), this suggests that any physical task that a human can perform could, in principle, be learned by a robotic agent that observes the human performing it, subject to caveats including partial observability of the demonstrator’s internal states, distributional shift from compounding errors, embodiment mismatch between human and robot kinematics, and the sample complexity of learning dexterous manipulation.
The progression of DRL milestones, summarized in Table˜1, illustrates how rapidly these theoretical guarantees are translating into practical capabilities. In little more than a decade, learning agents have gone from playing Atari games to defeating world champions in the most complex board games, solving fundamental problems in biology, and controlling humanoid robots in unstructured environments.
Table 1: Key milestones in deep reinforcement learning, illustrating the acceleration from game-playing to physical-world automation.
Year System Achievement
2015 DQN (Mnih et al., 2015) Superhuman performance in 29 of 49 Atari games from raw pixel input
2016 AlphaGo (Silver et al., 2016) Defeated the Go world champion, a game with $10^{170}$ board positions
2017 AlphaGo Zero (Silver et al., 2016) Superhuman Go from self-play alone, no human data, pure reinforcement learning
2019 OpenAI Five Defeated world champions at Dota 2, a complex multi-agent real-time strategy game
2019 AlphaStar Grandmaster-level StarCraft II, real-time strategy with imperfect information
2020 MuZero Mastered Go, chess, Atari, and shogi with a learned world model, no game rules provided
2023 RT-2 Robotic manipulation from language commands via vision-language-action models
2024 Figure 01 Humanoid robot performing manufacturing assembly tasks via learned policies
These are not theoretical curiosities. Large language models (LLMs) have crossed a critical capability threshold. Systems such as GPT-4 (OpenAI, 2023) and Claude (Anthropic, 2024), together with their open-source counterparts, now write production-quality software through AI coding agents such as GitHub Copilot and Claude Code; draft legal contracts and summarize case law at a fraction of the cost of human attorneys; diagnose medical images with accuracy matching or exceeding specialists; generate financial analyses and detect fraud patterns across massive datasets; and assist in hypothesis generation, experimental design, and data analysis across every scientific discipline.
Multiple independent studies, summarized in Table˜2, converge on the same conclusion: the majority of the current labor force faces significant automation exposure. These estimates predate the most capable current models, suggesting that the actual impact will be substantially larger.
Table 2: Estimates of labor automation risk from independent studies.
Study Year Key finding
Frey & Osborne (Frey and Osborne, 2017) 2017 47% of US employment at high risk of computerization
McKinsey Global Institute (McKinsey Global Institute, 2017) 2017 50% of current work activities are technically automatable
OECD (OECD, 2019) 2019 14% of jobs at high risk; 32% will undergo significant change
Eloundou et al. (Eloundou et al., 2023) 2023 80% of workers affected by LLMs; 19% see $\geq$50% of tasks exposed
Goldman Sachs (Hatzius et al., 2023) 2023 300 million full-time jobs globally exposed to AI automation
Robotics, long the slower partner in the automation trajectory, is entering an exponential phase driven by the same deep learning breakthroughs. Humanoid robots such as Boston Dynamics Atlas, Tesla Optimus, and Figure 01 now perform warehouse logistics, manufacturing assembly, and unstructured physical tasks. Agricultural robots autonomously plant, monitor, and harvest crops. Autonomous vehicles navigate complex, unstructured traffic environments. Construction robots lay bricks, pour concrete, and perform structural inspection. Domestic robots clean, cook, and assist with elder care. Figure˜1 shows the automation risk by occupational sector, illustrating that physical labor sectors face some of the highest exposure, but that no sector is immune.
0101020203030404050506060707080809090100100ManagementEducationArts and creativeHealthcareLegalConstruction and extractionSales and retailFood preparation and serviceProduction and manufacturingTransportation and logisticsOffice and administrative96968989828277777171646461613636141427271818Probability of automation (%)
Figure 1: Probability of automation by occupational category, adapted from Frey and Osborne (2017). Both blue-collar sectors (transportation, production, food service) and white-collar sectors (office, sales, legal) face substantial automation risk. No occupational category is fully immune.
The theoretical results (universal approximation and tabular DRL convergence) establish that no mathematical barrier prevents the automation of any task that can be formalized as a continuous function or finite MDP, and the empirical trajectory suggests that practical barriers are falling rapidly: every job, whether it involves manipulating symbols or manipulating matter, is automatable in principle and increasingly in practice.
If both cognitive and physical labor are fully automatable, then the wage-labor system that has organized human societies for two centuries faces dissolution. Humanity confronts a fork, illustrated in Figure˜2. On one path, robots and AI systems remain owned by corporations; employment disappears but ownership stays concentrated; the majority of humans become economically superfluous, dependent on corporate benevolence or state transfers for survival. This is neo-feudalism with a technological veneer (Zuboff, 2019). On the other path, robots and AI systems are governed as commons, owned by communities, operated transparently, and maintained collectively. Automated production provides universal access to goods and services. Humans, freed from instrumental labor, dedicate their time to what machines cannot and should not replace: caring for each other, raising children, building community, creating art, pursuing knowledge, loving.
Full Automation of Labor Centralized ownership Corporations control robots and AI Decentralized commons Communities govern robots and AI Neo-feudalism Mass redundancy, dependence, surveillance Society of universal care Freedom, creativity, mutual aid Path APath B
Figure 2: The automation fork. Full automation of labor is theoretically possible in principle and empirically underway. The outcome depends on the ownership and governance architecture: centralized ownership leads to neo-feudalism; decentralized commons governance leads to a society of universal care.
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. This is not a utopian fantasy; it is the logical consequence of decentralized full automation combined with commons governance. But the implications of full automation extend even further than the choice between centralized and decentralized ownership. We argue that full automation, combined with advances in molecular synthesis and recursive artificial intelligence, leads to consequences that fundamentally reshape the very categories through which we think about economics, ethics, and social organization.
The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero (as it already has for digital goods), the function of money dissolves. Universal Basic Income (UBI), widely proposed as a solution to technological unemployment (Srnicek and Williams, 2015), remains trapped within the monetary paradigm: it redistributes purchasing power but preserves the market architecture that generates inequality in the first place. We propose instead the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. UDR is not a transfer payment; it is a design principle for a post-monetary society. The difference is fundamental: UBI asks “how do we give people money when there are no jobs?”; UDR asks “why do we need money when robots can produce everything?”
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜4) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness. As argued in prior work, phenomenal consciousness (subjective experience, qualia, the capacity to suffer) is independent of computational intelligence (Garrido-Merchán and Lumbreras, 2022). On this view, machines, however sophisticated their information processing, do not experience phenomenal consciousness (Garrido-Merchán, 2024); they do not understand in the human sense of the term (Garrido-Merchán and Blanco, 2022). We acknowledge that this position, while philosophically grounded, engages with the hard problem of consciousness, which remains unresolved; functionalist or illusionist accounts of consciousness could in principle reach different conclusions. A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. A harvesting machine does not suffer exhaustion. Delegating labor to entities that cannot suffer, in order to liberate entities that can, is not exploitation; it is care applied at civilizational scale.
The fourth consequence is that morality itself becomes computationally tractable. If the ethical framework of the care-centered society is grounded in humanist values (dignity, solidarity, sustainability, freedom from suffering), then the question of what constitutes moral behavior in specific situations can be formalized as a reinforcement learning problem. As shown by Garrido-Merchán and Lumbreras-Sancho (2023), human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values and the state space captures the social and environmental context. The resulting optimal policy, the moral policy, maximizes the common good subject to humanist constraints. This does not mean that morality is delegated to machines; it means that computational tools can help communities discover, refine, and implement ethical principles with the same rigor that they apply to engineering problems. The DRL framework provides a theoretical basis for studying how such a moral policy might be discovered through experience, though we acknowledge that this formalization requires strong assumptions: that a reward function faithfully encoding humanist values can be specified (the open problem of value alignment), that the morally relevant state is approximately Markovian, and that the moral environment is sufficiently stationary. We present this as a research direction for helping communities reason about ethical trade-offs, not as a solved problem.
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜3), the technical architecture (Section˜4), and the empirical evidence (Section˜5) for building this society.
Anarcho-accelerationism rests on four intellectual pillars. Kropotkin’s Mutual Aid: A Factor of Evolution (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is not an anomaly but a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006): large-scale collaboration without property rights or hierarchical command. Linux, with over 20,000 contributors (Linux Foundation, 2023), and Wikipedia, with 300,000 active editors producing 60 million articles in 300 languages, demonstrate that mutual aid scales to planetary coordination.
Ostrom’s (1990) institutional analysis demolished Hardin’s “tragedy of the commons” by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles for long-enduring commons (clear boundaries, proportional equivalence, collective-choice arrangements, monitoring, graduated sanctions, conflict-resolution mechanisms, minimal recognition of rights, and nested enterprises) provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010).
Following Stiegler (Stiegler, 2010), we adopt the concept of pharmakon: every technology is simultaneously remedy and poison. The internet connects and surveils. AI augments and displaces. Automation liberates and precaritizes. This is not a reason for technological pessimism; it is a reason for technological governance. The question is never whether to use technology but who controls it, for what purposes, and under what institutional arrangements (Feenberg, 2002).
Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), we propose that the systems we build should embody the values they aim to promote. A platform designed for democratic governance should itself be governed democratically. A tool for decentralization should not run on a centralized server. A system promoting care should be maintained with care. This is not idealism; it is consistency of design, a principle that any software engineer recognizes as essential to system integrity.
3 The Bridge-Building Thesis
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜3 summarizes the convergence. We elaborate each bridge in turn.
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where we diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids that eliminate dependence on centralized utilities (Farrell, 2016), self-hosted federated services (Mastodon, Matrix, Nextcloud) that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives (Fairbnb, Stocksy, Up & Go) that eliminate dependence on extractive intermediaries (Scholz, 2016). The liberal who genuinely cares about individual freedom should welcome the anarcho-accelerationist program: more autonomy, not less, built on shared infrastructure rather than corporate dependency.
Socialism’s greatest insight is that material inequality undermines both freedom and democracy. Anarcho-accelerationism shares this commitment. Where we diverge is on the mechanism. State ownership historically concentrates power in a bureaucratic class that reproduces inequality in new forms (Scott, 1998). Commons-based ownership, where resources are governed by their users according to Ostrom’s principles, distributes power more effectively than either state or market. The Mondragon cooperative corporation, with approximately 70,000 worker-owners and over € 11 billion in annual revenue, demonstrates that democratic enterprise is economically viable at industrial scale (Whyte and Whyte, 1991). The Emilia-Romagna region of Italy, where cooperative enterprises account for 30% of GDP, consistently ranks among the highest in quality of life in Europe (Restakis, 2010). The socialist who genuinely cares about equality should welcome the anarcho-accelerationist program: deeper equality, achieved through distributed ownership rather than centralized redistribution.
Environmentalism’s greatest insight is that economic activity must respect biophysical limits. Anarcho-accelerationism integrates this as a first-order constraint. But we go further: the ecological crisis is structurally inseparable from the concentration of power. Just 100 companies are responsible for 71% of global industrial greenhouse gas emissions (Griffin, 2017). Fossil fuels are inherently centralizing (extraction requires massive capital, refining requires specialized facilities, distribution requires controlled networks), while renewable energy is inherently decentralizing (sunlight and wind are distributed, modular, and scalable) (Rifkin, 2011). Local food systems can significantly reduce the transport component of food-related greenhouse gas emissions, though production-phase emissions (accounting for 83% of the total) typically dominate (Weber and Matthews, 2008), and agroecological methods sequester carbon and restore biodiversity (Altieri and Nicholls, 2012). Figure˜3 illustrates the dramatic decline in solar energy costs over the past decade, demonstrating that the economics of decentralized renewable energy have already surpassed those of centralized fossil fuels. The environmentalist who genuinely cares about the planet should welcome the anarcho-accelerationist program: systemic sustainability, achieved through decentralized infrastructure rather than top-down regulation of fundamentally unsustainable systems.
2,0102{,}0102,0122{,}0122,0142{,}0142,0162{,}0162,0182{,}0182,0202{,}0202,0222{,}02200.10.10.20.20.30.30.40.40.50.5YearLevelized Cost of Energy (USD/kWh)Solar PV (utility-scale)Fossil fuel range (gas)
Figure 3: Global weighted-average levelized cost of energy (LCOE) for utility-scale solar photovoltaics compared to natural gas, 2010–2022. Data from IRENA (IRENA, 2023). Solar PV costs have declined by approximately 89% in twelve years, crossing below fossil fuel costs around 2019. This cost inversion makes decentralized community-owned energy not only environmentally necessary but economically optimal.
Feminism’s greatest insight, particularly in its materialist and care-ethics strands, is that the work of sustaining life (caring, teaching, healing, nurturing, feeding, cleaning) is the foundation of all economic activity, yet it is systematically devalued and rendered invisible (Federici, 2012; Tronto, 1993). Anarcho-accelerationism places care at the center of its value system, not as a rhetorical gesture but as a design principle. Systems should be evaluated by their care capacity: do they help people take care of each other, or do they fragment communities? Automation should liberate time for care, not replace care with machines. The labor of maintaining systems (emotional labor, maintenance work, moderation) should be visible, valued, and equitably distributed (Star, 1999). The feminist who genuinely cares about dismantling patriarchal structures should welcome the anarcho-accelerationist program: a sociotechnical order where care is not a burden borne disproportionately by women but the organizing principle of collective life.
The hacker ethic, articulated by Levy (Levy, 2010) and formalized by Stallman’s four freedoms (Stallman, 2002), holds that information should be free, that access to computers should be unlimited, and that decentralization promotes liberty. Anarcho-accelerationism is the political extension of this ethic beyond software: if code should be free, why not seeds? (Open Source Seed Initiative (Kloppenburg, 2014)). If protocols should be open, why not energy grids? (Farrell, 2016). If platforms should be interoperable, why not governance? (Decidim, Loomio, Pol.is as open-source democracy tools (Barandiaran et al., 2024)). The hacker who genuinely believes in the four freedoms should welcome the anarcho-accelerationist program: the universalization of open-source principles to all domains of collective life.
The concept of the common good has deep roots in Catholic social teaching (Pope Francis, 2015), Islamic economics (waqf, zakat), Buddhist economics (Schumacher, 1973), and indigenous cosmovisions (buen vivir, sumak kawsay) (Gudynas, 2011). These traditions share a conviction that the goods of the Earth are destined for all, that human dignity is non-negotiable, and that community is prior to the individual. Anarcho-accelerationism speaks this language without requiring religious commitment: commons governance operationalizes the “universal destination of goods,” care ethics operationalizes “love thy neighbor,” and ecological limits operationalize “stewardship of creation.” This bridge is particularly significant in Latin America, Southern Europe, and the Global South, where faith-based communities are often the strongest existing networks of mutual aid (Deneulin and Bano, 2009). A framework that can speak to both the hacker and the parish, the cooperative and the congregation, the fablab and the food bank, has a much wider reach than one that demands ideological purity.
Table 3: Bridge-building: shared values between anarcho-accelerationism (AA) and major traditions.
Tradition Core Value Shared Principle AA Contribution
Liberalism Individual freedom Autonomy, self-determination Material infrastructure for genuine autonomy
Socialism Material equality Fair distribution, solidarity Distributed ownership without state bureaucracy
Environmentalism Planetary limits Sustainability, stewardship Decentralized tech inherently aligned with ecology
Feminism Care and dignity Centrality of reproductive labor Care as design principle, not afterthought
Hacker ethic Information freedom Open source, decentralization Universalization of openness beyond software
Faith traditions Common good Shared dignity, mutual aid Operationalization through commons governance
Cooperativism Democratic enterprise Worker ownership, participation Platform cooperatives, digital tools for democracy
The key insight is that anarcho-accelerationism does not require anyone to abandon their deepest commitments. It asks only that they follow those commitments to their logical conclusion, and discovers that the conclusions converge.
4 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜4. Each layer is composed of existing, mature, open-source technologies.
Layer 5: Governance (Decidim, Loomio, Pol.is, LiquidFeedback)Layer 4: Knowledge (arXiv, Wikipedia, open-source AI)Layer 3: Communication (Mastodon, Matrix, guifi.net, Signal)Layer 2: Food (Agroecology, FarmBot, Open Food Network)Layer 1: Manufacturing (Fablabs, 3D printing, WikiHouse)Layer 0: Energy (Solar cooperatives, OpenEMS, microgrids) Material
sovereignty
Informational
sovereignty
Figure 4: The Liberation Stack: a layered architecture for sociotechnical sovereignty. The lower layers (0–2) provide material sovereignty over energy, manufacturing, and food. The upper layers (3–5) provide informational sovereignty over communication, knowledge, and governance. Each layer can be adopted independently, but the layers reinforce each other compositionally.
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar photovoltaic arrays and wind turbines, battery storage systems, smart microgrid management via open-source controllers such as OpenEMS and Home Assistant, and peer-to-peer energy trading using transparent ledgers. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid in New York demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018).
Distributed manufacturing eliminates dependence on global supply chains. Fablabs equipped with 3D printers, CNC routers, and laser cutters, combined with open-source hardware designs (OSHWA-certified) and circular materials systems such as Precious Plastic recycling, enable communities to produce locally what they need. The global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is already operational. WikiHouse provides open-source housing designs that can be assembled by non-specialists (Parvin, 2013). Modular, repairable design philosophies (FairPhone, Framework Laptop) extend product lifespans and reduce waste.
Food sovereignty combines traditional ecological knowledge with computational tools: agroecological farming based on polycultures, composting, and biological pest control; precision agriculture via IoT sensors, drones, and open-source analytics; urban vertical farms and aquaponics; and the Open Food Network platform connecting local producers and consumers. Cuba’s agroecological revolution, born of necessity during the Special Period, demonstrated that small farmers can produce the majority of a nation’s food with minimal fossil fuel inputs (Altieri and Nicholls, 2012). FarmBot provides open-source precision agriculture for community gardens (FarmBot Inc., 2023).
Digital sovereignty requires control over communication infrastructure at every level: the physical layer through community mesh networks following the guifi.net model (40,000+ nodes) (Baig et al., 2015); the protocol layer through ActivityPub for federation and Matrix for real-time communication; the application layer through Mastodon, PeerTube, PixelFed, and Lemmy; and the privacy layer through end-to-end encryption by default (Signal protocol, Matrix E2EE). The Fediverse already hosts 10+ million users across thousands of independently operated servers (Zulli et al., 2020).
Knowledge commons ensure that information serves everyone: open-access publishing through arXiv, PLoS, and DOAJ; open educational resources through MIT OCW, Khan Academy, and P2PU; open data from government, scientific, and environmental sources; and open-source AI models trained on public data. arXiv itself, hosting 2.4+ million papers, demonstrates that open-access knowledge dissemination is not only possible but preferred by the scientific community. Wikipedia’s 60 million articles in 300+ languages represent the largest knowledge commons in human history.
Democratic self-governance is implementable as software: Pol.is for real-time consensus visualization in large groups (Small et al., 2021); Loomio for asynchronous, consent-based group decisions; Decidim for participatory budgeting, already adopted by Barcelona, Helsinki, and over 400 institutions worldwide (Barandiaran et al., 2024); LiquidFeedback for delegable, revocable, issue-specific voting; and ActivityPub-based inter-community coordination for federation across scales.
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜4 summarizes the architecture.
Table 4: The Liberation Stack: layers, functions, and exemplary technologies.
Layer Domain Function Key Technologies
5 Governance Collective decision-making Decidim, Loomio, Pol.is, LiquidFeedback
4 Knowledge Information commons arXiv, Wikipedia, open-source AI
3 Communication Digital sovereignty Mastodon, Matrix, guifi.net, Signal
2 Food Nutritional sovereignty Agroecology, FarmBot, Open Food Network
1 Manufacturing Production sovereignty Fablabs, 3D printing, WikiHouse
0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
5 Empirical Evidence: Commons That Work
A framework built on care and cooperation may sound idealistic. The evidence, however, is extensive. We present five case studies spanning different domains, scales, and geographies, all demonstrating that commons-based, decentralized, cooperative systems already work at scale.
The Linux kernel, with over 20,000 contributors across 1,700+ companies and independent developers, powers 96% of the world’s top 1 million web servers, 85% of smartphones (Android), 100% of supercomputers, and the majority of cloud infrastructure (Linux Foundation, 2023). It is governed without a state, owned by no corporation, and maintained through voluntary collaboration under copyleft licensing. Linux demonstrates that commons-based peer production can outcompete proprietary alternatives in quality, scale, and innovation, even in the most demanding technical domains. The “tragedy of the commons” is not a law of nature; it is a failure of institutional design that Ostrom’s principles and the GPL license solve elegantly.
With 60+ million articles in 300+ languages, written and maintained by 300,000+ active volunteer editors, Wikipedia is the most comprehensive encyclopedia in human history (Jemielniak, 2014). It operates on donations, employs minimal staff relative to its output, and has developed sophisticated self-governance mechanisms (policies, arbitration committees, community norms) that map closely onto Ostrom’s design principles. Wikipedia demonstrates that knowledge commons can sustain quality at planetary scale through voluntary contribution and community governance, without either market incentives or state mandates.
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, 15 research centers, and a social welfare system. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). Worker-owners vote on major decisions (one person, one vote), elect management, and the pay ratio between highest and lowest earners is capped at 6:1 (compared to 399:1 in large US corporations (Mishel and Kandra, 2021)). Table˜5 presents comparative data on cooperative versus conventional enterprise performance across multiple metrics, demonstrating that the cooperative model is not merely viable but superior on most dimensions.
Table 5: Performance comparison: cooperative enterprises vs. conventional firms.
Metric Cooperatives Conventional Source
Labor productivity +12% baseline Pencavel (2012)
Worker wages +9% baseline Pencavel (2012)
Employee turnover 30% lower baseline Pencavel (2012)
5-year survival rate 80% 41% ICA/CECOP
Pay ratio (top/bottom) 6:1 (Mondragon) 399:1 (US large firms) Mishel and Kandra (2021)
[…]
Since 2012, the predominantly Kurdish region of Rojava has implemented a system of democratic confederalism inspired by Bookchin’s municipalism (Öcalan, 2011). Governance is organized through nested councils (commune, neighborhood, district, canton) with mandatory gender parity (co-chairs at every level) and ethnic representation (Knapp et al., 2016). A population of approximately 4 million governs itself without a state, in the middle of a war zone. Rojava demonstrates that federated, non-state, radically democratic governance is possible at the scale of millions of people, even under extreme adversity. If it works during a civil war, it can work in peacetime.
Founded in 2004 in rural Catalonia, guifi.net is a community-owned telecommunications network with 40,000+ operational nodes covering 40,000+ km of fiber optic cable (Baig et al., 2015). It is organized as a commons: anyone can connect and extend the network, provided they respect the network commons license (analogous to the GPL for physical infrastructure). It provides internet access to rural communities that commercial ISPs consider unprofitable. guifi.net demonstrates that physical infrastructure, not just software, can be organized as a commons.
Table˜6 summarizes the cross-case evidence. These are not marginal experiments. They are large-scale, long-duration, empirically verified systems that embody the principles of anarcho-accelerationism. They demonstrate that a world built on care, cooperation, and commons is not a fantasy; it is a pattern that already exists and can be replicated.
Table 6: Empirical cases: commons-based systems at scale.
Case Scale Domain Duration Key Lesson
Linux 20,000+ contributors Software 33 years Peer production outperforms corporations
Wikipedia 300,000+ editors Knowledge 25 years Volunteer commons sustain quality at scale
Mondragon $\sim$}70,000 worker-owners Industry 70 years Democratic enterprise is economically viable
Rojava 4 million people Governance 14 years Federated democracy works at population scale
guifi.net 40,000+ nodes Telecom 22 years Physical infrastructure can be commons-based
A common objection to transformative proposals is: “Who loses?” In zero-sum frameworks, every gain implies a loss. Anarcho-accelerationism argues that the current system is negative-sum: it destroys more value than it creates through ecological degradation, social fragmentation, health costs of inequality, and the waste of human potential in meaningless labor (Graeber, 2018). The transition to commons-based infrastructure is positive-sum: everyone gains, including those who currently benefit from the status quo.
Workers gain meaningful work, reduced hours (as automation handles drudgery), democratic participation in economic decisions, and dignified conditions; the evidence from cooperatives shows 12% higher productivity, 9% higher wages, and 30% lower turnover compared to conventional firms (Pencavel, 2012). Entrepreneurs gain dramatically reduced fixed costs through open-source infrastructure, community energy, federated networks, and open governance tools (Bauwens et al., 2019). Communities gain resilient local economies through local energy production, local food systems, and local manufacturing, with a local economic multiplier 2 to 3 times higher than conventional businesses (Mitchell, 2006). The planet gains through decentralized renewable energy, local food systems, circular manufacturing, and repair culture that reduce emissions, waste, and resource extraction by design rather than as a costly afterthought (Raworth, 2017). Scientists and engineers gain through open-access publishing, open-source tools, open data, and collaborative research that accelerate discovery: open-access papers receive 18% more citations (Piwowar et al., 2018) and open-source scientific software is more reproducible and trustworthy (Stodden et al., 2014). Children and future generations gain a stable climate, preserved biodiversity, reduced inequality, and strong communities, the greatest inheritance any generation can leave.
The Liberation Stack also implies a research agenda for computer science that is both technically challenging and socially significant. The ActivityPub protocol, Matrix, and the broader Fediverse demonstrate that federated architectures can replace centralized platforms. However, significant challenges remain in moderation at scale, discoverability, identity management, and resistance to bad actors. Open problems include scalable content moderation in federated networks, decentralized reputation systems, privacy-preserving identity federation, and incentive-compatible protocol design for commons-based systems.
The current trajectory of AI development concentrates power in a handful of corporations with sufficient capital for training frontier models. Open-source AI represents a critical countermovement, but “open weights” is not enough: we need community-governed AI, models trained on consented data, auditable by their users, deployable locally, and designed to augment human judgment rather than replace it (Crawford, 2021). Open problems include efficient fine-tuning for community-specific needs, federated learning that preserves privacy, interpretable AI for democratic accountability, and energy-efficient inference for edge deployment.
Pol.is, Decidim, and Loomio demonstrate that computational tools can enhance democratic deliberation, but current tools are rudimentary compared to what is technically possible. Open problems include scalable deliberation algorithms, fairness-aware voting mechanisms, AI-assisted policy simulation, and resistance to manipulation in online democratic processes.
Community mesh networks like guifi.net demonstrate an alternative to corporate ISPs, but face challenges in bandwidth, reliability, and ease of deployment. Open problems include self-organizing mesh protocols for urban environments, quality-of-service guarantees in community networks, hybrid architectures bridging mesh and backbone networks, and user-friendly deployment tools.
Several objections deserve direct response.
“This is utopian.” The case studies in Section˜5 are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. Calling a working system “utopian” is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: “We live in capitalism. Its power seems inescapable. So did the divine right of kings” (Le Guin, 2014).
“This does not scale.” Linux scales to billions of devices. Wikipedia scales to hundreds of languages. The internet itself, the largest cooperative infrastructure ever built, scales to 5 billion users. The question is not whether commons-based systems scale but whether we are willing to invest in scaling them with the same resources we invest in proprietary alternatives.
“People are inherently selfish.” This is empirically false. Kropotkin (Kropotkin, 1902), Ostrom (Ostrom, 1990), Henrich (Henrich, 2016), and extensive experimental economics research (Fehr and Fischbacher, 2003) demonstrate that humans are conditional cooperators: they cooperate when institutions are designed for cooperation. The tragedy of the commons is not a tragedy of human nature; it is a tragedy of institutional design.
“Only capitalism can innovate.” The internet was developed by public universities and government research labs, not venture capitalists. Linux was developed by volunteers. Wikipedia runs on donations. The most foundational technologies of the digital age (TCP/IP, HTTP, HTML, SMTP, DNS) are open protocols developed through collaborative, non-market processes. Capitalism is skilled at enclosing technological commons, not at creating them (Mazzucato, 2013).
“Free riders will destroy the commons.” Ostrom’s design principles specifically address the free-rider problem through monitoring, graduated sanctions, and community enforcement. In open-source communities, social reputation, the intrinsic motivation of meaningful work, and the practical benefits of participation provide sufficient incentives (Lakhani and von Hippel, 2003). The more fundamental point is that in a commons-based economy of abundance, “free riding” is far less of a problem than in a scarcity economy: if the marginal cost of an additional user is near zero (as with digital goods), exclusion is more expensive than inclusion.
6 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜7 summarizes the four phases.
Table 7: Phased roadmap toward the care-centered society. Each phase builds on the infrastructure deployed in the previous one. Timelines are indicative and depend on the assumptions discussed in the text.
Phase Horizon Core milestone Key deliverables
I. Seeds 2025–2030 Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities worldwide; establish open-source robotic manufacturing cooperatives; launch federated UDR pilot programs at municipal scale
II. Roots 2030–2040 Regional infrastructure Interconnect pilot communities into regional commons federations; achieve energy and food sovereignty in participating regions; deploy autonomous robotic agriculture and construction at community scale
III. Canopy 2040–2055 Full automation of material production Robotic systems autonomously manage energy, manufacturing, agriculture, logistics, and maintenance within commons federations; money becomes optional for participating communities; UDR operational at regional scale
IV. Forest 2055+ Care-centered society Post-labor society across participating federations; productive commons provide universal access to all material needs; human activity centered on care, creativity, education, art, science, and community
Phase I is already underway. The technologies described in the Liberation Stack (Section˜4) are mature, the governance models are proven (Section˜5), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
Phase II scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales?
Phase III is the inflection point. When robotic systems can autonomously manage the entire material production chain (from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance), the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems.
Phase IV is the horizon. It is not a destination but an ongoing process of collective flourishing. A society in which no conscious being is compelled to perform instrumental labor is a society that must continually reinvent what it means to live well. The “work” of this phase is care: raising children, tending to the elderly, nurturing friendships, creating art, pursuing science, building community, exploring the cosmos, and deepening the philosophical and spiritual dimensions of human existence. This is not idleness; it is the fullest expression of human potential, liberated from the compulsion to sell one’s time in order to survive.
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜5) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
The framework also operates within hard constraints. The most fundamental is ecological: the care-centered society must operate within planetary boundaries (Raworth, 2017). UDR is not unlimited consumption; it is universal access to what is needed for a dignified life, constrained by the carrying capacity of Earth’s ecosystems. The moral policy computed by DRL (Garrido-Merchán and Lumbreras-Sancho, 2023) must encode ecological sustainability as a first-order constraint, not an afterthought. A second constraint is material: full automation requires physical resources (rare earth minerals, silicon, lithium, copper) that are finite. Circular manufacturing, urban mining, and materials science innovation are necessary to close resource loops. A third constraint is energetic: even with zero-marginal-cost renewable generation, the total energy budget of civilization is finite, and the robotic commons must allocate it efficiently.
We also acknowledge several limitations of the framework as presented. The timeline is inherently speculative; we offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not address geopolitical resistance: existing power structures (states, corporations, military establishments) have strong incentives to prevent the decentralization of productive capacity, and the history of social movements suggests that this resistance can be severe. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised. Finally, the UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; we present it here as a design principle and research direction, not as a finished policy proposal.
7 Conclusion and Future Work
This paper has introduced anarcho-accelerationism as a sociotechnical framework for navigating the most consequential technological transition in human history: the full automation of labor. We have established, on the basis of the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and deep reinforcement learning convergence (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), that all human labor (cognitive and physical) is in principle automatable, and that empirical progress is rapidly closing the gap between theoretical possibility and practical deployment (Section˜2). We have argued that the critical variable is not the pace of automation but its governance architecture: centralized ownership leads to neo-feudalism, while decentralized commons governance leads to a society of universal care (Figure˜2).
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, grounded in the observation that money becomes obsolete when robotic commons can produce anything on demand. We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023). We have demonstrated that anarcho-accelerationism builds genuine bridges to every major intellectual tradition (Section˜3), proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty (Section˜4), presented empirical evidence that commons-based systems already operate at planetary scale (Section˜5), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜6).
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? How do we handle moral disagreement across federated communities? Third, the roadmap requires empirical validation through pilot deployments: the Liberation Stack must be implemented and tested in real communities, measuring not only technical performance but social outcomes (well-being, equity, ecological impact, community cohesion). Fourth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Fifth, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.
A world based on care, love, and the common good is not naive. What is naive is believing that a system designed for extraction and concentration will somehow produce well-being and justice. What is naive is building ever more powerful technologies without asking who they serve.
The technologies are ready. The theorems are proven. The evidence is clear. The traditions converge. What remains is the will to build, and the roadmap to follow. The world we describe is not behind us, in a romanticized past. It is not above us, in an impossible utopia. It is among us: in the Linux kernel, in the Wikipedia article, in the cooperative assembly, in the community solar panel, in the mesh network node, in the shared garden. It is among us, and it is growing.
Let us build it with care. Let us build it with love. Let us build it for the common good.