### v6

# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

Abstract
Foundational results in machine learning establish that all human labor may in principle be automatable. Without deliberate intervention, this trajectory risks concentrating productive capacity in a handful of corporations, resulting in techno-feudalism: mass economic redundancy, surveillance-based control, and dependence on corporate benevolence for survival. To avert this outcome, this paper introduces peaceful anarcho-accelerationism, a rigorously defined sociotechnical framework grounded in the 200-year anarchist tradition from Godwin through Kropotkin to Bookchin, and in the methodological categories of Eltzbacher, Nettlau, and Corrêa, for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. Specifically, I state five formal hypotheses and six research objectives, present a formal definition through analytical categories of interdependent spheres, and propose the Liberation Stack as a layered technical architecture with explicit preconditions and gate conditions for each layer. Furthermore, I introduce Universal Desired Resources as a post-monetary design principle that eliminates the material basis of intersectional oppression, and develop a framework for progressive state dissolution through incremental, reversible commons-building compatible with existing democratic institutions. Additionally, a nonviolent social mobilization strategy maps concrete peaceful methods to each stage of transition. Moreover, I show that accelerationism and degrowth share anarchist pacifism as substrate and differ only along a Pareto-optimal technological frontier. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary commons initiatives confirms that commons-based systems already operate at scale. Finally, I conclude with a phased roadmap specifying explicit assumptions, hard constraints, gate conditions between phases, and detailed limitations.
Keywords: anarcho-accelerationism; full automation; commons governance; universal desired resources; accelerationist ecologism; nonviolent strategy; Fediverse
1 The Convergence Window and the Automation Imperative
We live in a paradox. Humanity possesses the technical capacity to feed, house, heal, educate, and connect every person on Earth (Bookchin, 1971). Solar energy is now the cheapest electricity source in history (IRENA, 2023). A single modern smartphone carries more computational power than the Apollo mission control. Open-source software runs 96% of the world’s top web servers (Linux Foundation, 2023). And yet 735 million people face chronic hunger (FAO et al., 2023), three individuals hold more wealth than the bottom half of the United States (Oxfam International, 2023), and the six largest technology corporations exercise more influence over global communication than any state in history (Zuboff, 2019).
This disparity suggests that the central challenge is not one of resource scarcity but of sociotechnical architecture: the systems governing production, distribution, and access are structured to concentrate rather than distribute.
Before proceeding, I state the formal hypotheses that this paper investigates and the research objectives it pursues, so that the reader can evaluate the argument against explicit claims rather than vague aspirations.
The first hypothesis (H1) is that all human labor, both cognitive and physical, is in principle automatable. This claim rests on the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and the convergence results of deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), which together establish that no mathematical barrier prevents the automation of any formalizable task or any learnable sensorimotor policy. I emphasize that this is an in-principle claim: the existence of approximating networks and convergent policies does not guarantee tractable training, nor does it address tasks whose specification may be inherently incomplete. The second hypothesis (H2) is that decentralized commons governance of automation produces better outcomes than centralized corporate ownership, measured along three dimensions: inequality (Gini coefficient and wealth distribution), human well-being (life satisfaction, health outcomes, educational attainment, and Maslow-level satisfaction), and ecological footprint (carbon emissions, resource throughput, and biodiversity impact). This hypothesis is testable through comparative analysis of existing commons-based systems and through pilot implementations of the Liberation Stack. The third hypothesis (H3) is that nonviolent social mobilization can achieve the commons transition without state seizure, drawing on the empirical findings of Chenoweth and Stephan (2011) that nonviolent campaigns succeed at twice the rate of violent ones, and on the strategic framework of Sharp (1973) that identifies 198 methods of nonviolent action applicable to the commons transition. The fourth hypothesis (H4) is that the apparent opposition between degrowth and accelerationism is a false binary: both positions share anarchist pacifism as their ethical substrate and occupy different points along a Pareto-optimal frontier trading off ecological risk against human suffering during the transition. A mature movement would support both strategies simultaneously, and technological progress shifts the entire frontier toward the origin. The fifth hypothesis (H5) is that each layer of the Liberation Stack has identifiable preconditions for successful implementation, and that the layers form a partial ordering of dependencies: energy is the precondition for all material layers, communication is the precondition for knowledge and governance, and no layer can be sustainably deployed without the layers below it in the dependency graph.
The research objectives follow from these hypotheses. The first objective (O1) is to define peaceful anarcho-accelerationism rigorously through the analytical categories of Corrêa (Corrêa, 2015) and the comparative method of Eltzbacher (Eltzbacher, 1908), situating it within the 200-year anarchist tradition and demonstrating its continuity with classical principles. The second objective (O2) is to propose a concrete technical architecture, the Liberation Stack, with explicit preconditions and gate conditions for each layer, a dependency graph specifying the partial ordering of implementation, and a framework for progressive state dissolution through incremental commons-building. The third objective (O3) is to develop a nonviolent mobilization strategy grounded in the empirical evidence of Chenoweth and Stephan (2011) and the strategic repertoire of Sharp (1973), mapping specific nonviolent methods to each stage of the commons transition. The fourth objective (O4) is to demonstrate that intersectional liberation, as theorized by Crenshaw (1989) and Collins (2019), is a structural consequence of Universal Desired Resources: when the material basis of all axes of oppression is eliminated simultaneously, intersectional liberation follows as a design outcome rather than a political aspiration. The fifth objective (O5) is to provide empirical evidence from existing commons-based systems, including Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary initiatives such as the Kudumbashree network, the Preston Model, Cooperation Jackson, and the Evergreen Cooperatives, demonstrating that the organizational principles of the framework are not speculative but operational. The sixth objective (O6) is to present a phased roadmap with explicit assumptions, preconditions, gate conditions between phases, hard constraints, detailed limitations, and directions for future research, so that the framework can be evaluated, tested, and refined by the broader scholarly community.
The dominant architectures of our sociotechnical systems, namely centralized cloud platforms, extractive business models, proprietary codebases, fossil-fuel energy grids, and top-down governance, were designed to concentrate power. They are spectacularly effective at this task. But they are not the only architectures available to us. A growing body of evidence from computer science, institutional economics, cooperative studies, and political ecology suggests that decentralized, commons-based, federated architectures can achieve equal or superior performance in coordination, resilience, innovation, and human well-being (Benkler, 2006; Ostrom, 1990; Scholz, 2016).
This paper introduces anarcho-accelerationism: a framework for sociotechnical system design that takes this evidence seriously. The name is deliberately provocative, but the substance is rigorous. The prefix anarcho- refers not to chaos but to the 150-year tradition of anarchist political philosophy, from Kropotkin’s empirical study of mutual aid in nature (Kropotkin, 1902) to Bookchin’s ecological municipalism (Bookchin, 2005) to Ostrom’s Nobel Prize-winning work on commons governance (Ostrom, 1990). The core principle is self-organization without domination. The suffix -accelerationism refers not to reckless technological enthusiasm but to the critical insight that current productive forces, if redirected from accumulation toward liberation, are sufficient to abolish artificial scarcity and enable universal flourishing (Srnicek and Williams, 2015). The core principle is technological sufficiency for post-scarcity. The synthesis yields a design philosophy centered on three values: care, understood as the design of systems that sustain life, relationships, and communities rather than maximize engagement, extraction, or growth (Federici, 2012; Puig de la Bellacasa, 2017); love, understood not as sentimentality but as political commitment, the design choice to build systems that strengthen bonds between people rather than atomize them (hooks, 2000; Fromm, 1956); and the common good, the conviction that resources essential for human dignity, such as energy, food, knowledge, communication, and code, should be governed as commons, not enclosed as property (Ostrom, 1990; Hess and Ostrom, 2007). As Nettlau wrote in surveying the entire history of anarchism, the movement is unified by a belief in reason and progress, and by the pursuit of the maximum possible realization of the liberty and wellbeing of all (Nettlau, 1996).
The urgency of this framework stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. I note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. My claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
The convergence theorems for deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015) establish that an agent interacting with any environment describable as a Markov Decision Process can learn an optimal policy through observation of states, actions, and rewards. For any MDP $(S,A,P,R,\gamma)$ with finite state and action spaces, Q-learning converges to the optimal action-value function $Q^{*}$ given sufficient exploration (Watkins and Dayan, 1992). With deep neural networks as function approximators, Q-learning has achieved remarkable empirical success in continuous, high-dimensional state spaces, though formal convergence guarantees do not extend from the tabular case to function approximation in general. Combined with imitation learning, where policies are acquired by observing human demonstrations (Schaal, 1999), this suggests that any physical task that a human can perform could, in principle, be learned by a robotic agent, subject to caveats including partial observability, distributional shift, embodiment mismatch, and sample complexity. The progression of DRL milestones, summarized in Table˜1, illustrates how rapidly these theoretical guarantees are translating into practical capabilities.
Table 1: Key milestones in deep reinforcement learning, illustrating the acceleration from game-playing to physical-world automation.
Year System Achievement
2015 DQN (Mnih et al., 2015) Superhuman performance in 29 of 49 Atari games from raw pixel input
2016 AlphaGo (Silver et al., 2016) Defeated the Go world champion, a game with $10^{170}$ board positions
2017 AlphaGo Zero (Silver et al., 2016) Superhuman Go from self-play alone, no human data
2019 OpenAI Five Defeated world champions at Dota 2, a complex multi-agent real-time strategy game
2020 MuZero Mastered Go, chess, Atari, and shogi with a learned world model, no game rules provided
2023 RT-2 Robotic manipulation from language commands via vision-language-action models
2024 Figure 01 Humanoid robot performing manufacturing assembly tasks via learned policies
Large language models have crossed a critical capability threshold. Systems such as GPT-4 (OpenAI, 2023) and Claude (Anthropic, 2024), together with their open-source counterparts, now write production-quality software, draft legal contracts, diagnose medical images with accuracy matching specialists, generate financial analyses, and assist in hypothesis generation across every scientific discipline. Multiple independent studies converge on the same conclusion: the majority of the current labor force faces significant automation exposure, as summarized in Table˜2.
Table 2: Estimates of labor automation risk from independent studies.
Study Year Key finding
Frey & Osborne (Frey and Osborne, 2017) 2017 47% of US employment at high risk of computerization
McKinsey Global Institute (McKinsey Global Institute, 2017) 2017 50% of current work activities are technically automatable
OECD (OECD, 2019) 2019 14% of jobs at high risk; 32% will undergo significant change
Eloundou et al. (Eloundou et al., 2023) 2023 80% of workers affected by LLMs; 19% see $\geq$50% of tasks exposed
Goldman Sachs (Hatzius et al., 2023) 2023 300 million full-time jobs globally exposed to AI automation
Robotics is entering an exponential phase driven by the same deep learning breakthroughs. Humanoid robots such as Boston Dynamics Atlas, Tesla Optimus, and Figure 01 now perform warehouse logistics, manufacturing assembly, and unstructured physical tasks. Figure˜1 shows the automation risk by occupational sector.
0101020203030404050506060707080809090100100ManagementEducationArts and creativeHealthcareLegalConstruction and extractionSales and retailFood preparation and serviceProduction and manufacturingTransportation and logisticsOffice and administrative96968989828277777171646461613636141427271818Probability of automation (%)
Figure 1: Probability of automation by occupational category, adapted from Frey and Osborne (2017). Both blue-collar and white-collar sectors face substantial automation risk. No occupational category is fully immune.
If both cognitive and physical labor are fully automatable, then the wage-labor system that has organized human societies for two centuries faces dissolution. Humanity confronts a fork, illustrated in Figure˜2. On one path, robots and AI systems remain owned by corporations; employment disappears but ownership stays concentrated; the majority of humans become economically superfluous, dependent on corporate benevolence or state transfers for survival. This is neo-feudalism with a technological veneer (Zuboff, 2019). On the other path, robots and AI systems are governed as commons, owned by communities, operated transparently, and maintained collectively. Automated production provides universal access to goods and services. Humans, freed from instrumental labor, dedicate their time to what machines cannot and should not replace: caring for each other, raising children, building community, creating art, and pursuing knowledge.
Full Automation of Labor Centralized ownership Corporations control robots and AI Decentralized commons Communities govern robots and AI Neo-feudalism Mass redundancy, dependence, surveillance Society of universal care Freedom, creativity, mutual aid Path APath B
Figure 2: The automation fork. Full automation of labor is theoretically possible in principle and empirically underway. The outcome depends on the ownership and governance architecture.
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. The implications of full automation extend further than this binary choice. The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero, the function of money dissolves. I propose the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022). A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. Delegating labor to entities that cannot suffer, in order to liberate entities that can, constitutes a systematic application of care ethics to the organization of production. The fourth consequence is that morality itself becomes computationally tractable: human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values (Garrido-Merchán and Lumbreras-Sancho, 2023), though this formalization requires strong assumptions about value alignment and moral stationarity.
Previous attempts at decentralized, commons-based social organization illuminate both the promise and the perils of the present moment. The Paris Commune of 1871 demonstrated that workers could govern a major city through direct democracy, establishing free education, gender equality in wages, and cooperative workshops, but it lasted only 72 days before being crushed by military force (Ross, 2015). The Makhnovshchina in Ukraine from 1918 to 1921 organized agricultural communes across a territory of seven million people under conditions of civil war, demonstrating the viability of federated anarchist governance at regional scale, but was destroyed by Bolshevik betrayal after having resisted both the White Army and foreign intervention (Malet, 1982). The Kronstadt rebellion of 1921 saw the very sailors who had enabled the Russian Revolution demand the soviet democracy they had been promised, only to be massacred by the regime they had helped create (Avrich, 1970). The Spanish Revolution of 1936 achieved the most extensive implementation of anarchist principles in history: the CNT-FAI collectivized factories, farms, and public services across Catalonia and Aragon, with industrial production maintained and in some cases increased under worker self-management (Bolloten, 1991). Agricultural collectives in Aragon organized production, distribution, and social services for hundreds of thousands of people through assembly-based governance (Casanova, 2006). Yet the revolution was destroyed by the convergence of Francoist military assault, Stalinist repression from within the Republican camp, and the CNT’s fateful decision to join the Republican government, contradicting its own anti-statist principles.
These movements failed not because their organizational principles were flawed, but because they lacked three conditions that are now, for the first time in history, within reach. First, they had no technological infrastructure for decentralized coordination at scale: no internet, no mesh networks, no federated protocols, no real-time communication across territories. The CNT’s extraordinary organizational achievement of reaching 1.5 million members relied on physical meetings, printed newspapers, and postal correspondence, all vulnerable to disruption by military force. Today, federated communication protocols (ActivityPub, Matrix), community mesh networks (guifi.net), and open-source governance tools (Decidim, Loomio) provide the infrastructure for coordination that the historical movements did not have access to. Second, they had no energy independence. Every historical anarchist territory depended on external energy sources controlled by hostile states. Community-owned solar microgrids, enhanced geothermal systems, and iron-air batteries now make energy sovereignty technically and economically feasible at the community scale. Third, they had no productive automation. The collectives of 1936 still required human labor for every productive task, making them vulnerable to the economic disruptions of war. Autonomous robotic systems for agriculture, manufacturing, and logistics now make economic self-sufficiency feasible without requiring the full participation of every community member in manual labor.
The remainder of this paper provides the formal definition (Section˜2), the political bridges (Section˜3), the technical architecture and ecological framework (Section˜4), the nonviolent mobilization strategy (Section˜5), the empirical evidence (Section˜6), and the roadmap (Section˜7) for the framework proposed in this paper.
2 Theoretical Foundations: A Formal Definition of Peaceful Anarcho-Accelerationism
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. My methodology draws on six major historians of anarchism and updates their analytical categories for the technological conditions of the 21st century, following the principle of totality and interdependence articulated by Corrêa (Corrêa, 2015).
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to my approach. Max Nettlau (Nettlau, 1996), known as the Herodotus of Anarchy, produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie. Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association. George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism through the Enlightenment to contemporary movements.
Peaceful anarcho-accelerationism takes specific contributions from each of Eltzbacher’s seven thinkers, as summarized in Table˜3. From Godwin (Godwin, 1793) it takes reason as the instrument of social progress: the conviction that rational inquiry, not authority, should guide human organization, and that education and rational discourse are the primary mechanisms of social change. From Proudhon (Proudhon, 1840) it takes the critique of property as monopoly and the proposal of mutualism, the blueprint for commons-based economic organization, and the principle of federation as the alternative to both state centralism and market atomization. From Stirner (Stirner, 2014) it takes the insistence that liberation must serve the concrete individual, not an abstract collective, a corrective against the authoritarian temptation in all revolutionary movements. From Bakunin (Bakunin, 1990) it takes the fundamental anarchist thesis that no state apparatus, however well-intentioned, can serve as an instrument of liberation, which grounds the insistence on stateless, federated governance. From Kropotkin (Kropotkin, 1902) it takes the scientific grounding of cooperation as a factor of evolution, the ethics of mutual aid, and the vision of a post-scarcity society achieved through the rational application of technology. From Tucker (Tucker, 1893) it takes the libertarian commitment to eliminating structural monopolies, updated for the 21st century as the monopolies of platform data, proprietary AI, centralized energy, and intellectual property. From Tolstoy (Tolstoy, 1894) it takes the absolute commitment to peaceful means, the rejection of revolutionary violence, and the conviction that the means must be consistent with the ends.
Table 3: Eltzbacher’s seven thinkers and the specific contribution of each to peaceful anarcho-accelerationism.
Thinker Period Contribution to peaceful anarcho-accelerationism
Godwin 1756–1836 Reason as instrument of social progress; education and persuasion over coercion
Proudhon 1809–1865 Mutualism; federation as alternative to state and market; commons-based economy
Stirner 1806–1856 Individual sovereignty; liberation must serve the concrete person, not abstractions
Bakunin 1814–1876 Anti-statism; no state apparatus can serve as instrument of liberation
Kropotkin 1842–1921 Mutual aid as evolutionary factor; post-scarcity through rational technology
Tucker 1854–1939 Elimination of structural monopolies (updated: data, AI, energy, IP)
Tolstoy 1828–1910 Principled nonviolence; means must be consistent with ends
Following the methodology of Corrêa (Corrêa, 2015), who insists on the principle of totality and interdependence in analyzing anarchist movements, I define peaceful anarcho-accelerationism through seven analytical categories that are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole. The seven categories, compared with their classical formulations, are presented in Table˜4.
Table 4: Seven analytical categories of anarchism, updated from classical anarchism to peaceful anarcho-accelerationism, following Eltzbacher’s comparative method and Corrêa’s principle of totality.
Category Classical Anarchism Peaceful Anarcho-Accelerationism
Ideology Anti-statism, anti-capitalism, voluntary association Techno-anarchism: commons-governed automation for universal care; anti-domination extended to algorithmic and platform power
[…]
Social force The proletariat, the peasantry, artisan classes The precariat (Standing, 2011), digital commons builders, platform cooperativists, maker communities, open-source developers
Power Rejection of all coercive power; direct action Distributed power through technology: federated governance, mesh networks, community-owned AI; power with, not power over
Domination State and capital as twin pillars of domination Techno-feudalism (Varoufakis, 2023): platform monopolies, surveillance capitalism (Zuboff, 2019), algorithmic management, data extractivism
Self-management Workers’ councils, syndicates, communes, assemblies AI-assisted commons governance: Decidim, Pol.is, LiquidFeedback; robotic production managed by community protocols
Social classes Bourgeoisie vs. proletariat Cloud lords vs. cloud serfs; platform owners vs. platform-dependent precariat; data extractors vs. data producers
Corrêa’s methodology requires that these categories be understood through the principle of totality and interdependence across three registers. The first register concerns the unity of theory, practice, and history. Peaceful anarcho-accelerationism is simultaneously a theoretical framework (a formal analysis of domination and liberation under conditions of full automation), a practical program (the Liberation Stack, the UBI-to-UDR transition, the nonviolent mobilization strategy), and a historical continuation of the anarchist movement (from Godwin’s rational anarchism through the Spanish collectives to the Fediverse). No register is intelligible without the others: theory without practice is academic speculation, practice without theory is blind activism, and both without historical grounding repeat the errors of the past. The second register concerns the unity of thought and action. The intellectual analysis of domination (identifying techno-feudalism, surveillance capitalism, algorithmic management as the contemporary forms of the state-capital nexus) must be inseparable from the construction of alternatives (building community energy cooperatives, deploying federated communication platforms, establishing open-source manufacturing). In the language of Corrêa’s especifismo, the political level of theoretical analysis must be embedded within the social level of popular movements and concrete organizing. This is what Corrêa calls social insertion: the practice of embedding anarchist principles within existing social movements rather than standing apart from them in ideological purity. The third register concerns the unity of form and content. The organizational form of the movement, federated, horizontal, commons-based, must reflect its content, which is liberation, care, and non-domination. A movement that advocates decentralization cannot be organized hierarchically. A framework that promotes care cannot treat its participants instrumentally. This is the principle of prefigurative consistency (Graeber, 2013): the means must embody the ends.
Corrêa further identifies social vectors, the specific social forces that carry the potential for transformation, and critiques, the specific structures of domination against which the movement is directed. In my framework, the social vectors are the precariat (the growing class of workers in precarious, platform-mediated employment), the open-source development community (which already practices commons-based peer production at planetary scale), the cooperative movement (from Mondragon to platform cooperatives), the agroecological movement, the community energy movement, and the Fediverse community (which builds and maintains decentralized communication infrastructure). The critiques are directed against techno-feudalism (the concentration of digital productive capacity in a handful of platform corporations), surveillance capitalism (the extraction and monetization of behavioral data), algorithmic management (the replacement of human supervisors with opaque automated systems), fossil-fuel centralization (the structural dependence on concentrated energy sources that enables authoritarian control), and intellectual property enclosure (the artificial scarcification of knowledge, seeds, and code through patent and copyright regimes).
The update from classical to accelerationist anarchism is not a rupture with the tradition but its logical extension to new conditions of domination. Where Proudhon confronted industrial capital, we confront platform capital. Where Bakunin organized workers against the factory owner, we organize communities against the algorithm. Where Kropotkin documented mutual aid in nature, we document mutual aid in open-source software. The principle remains constant: self-organization without domination. The terrain has shifted from the factory floor to the digital commons.
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), I articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them. The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality. The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination (Ward, 1973).
Time1793184019021936194819731994200520122026 Godwin PJ Proudhon Property Kropotkin Mutual Aid Spanish Revolution Zapatistas EZLN Rojava Confederalism Anarchist Bookchin Post-Scarcity Stiegler Pharmakon Srnicek Williams Anarcho- Accel. Accelerat. Godwin Reason Tolstoy Nonviolence Gandhi Satyagraha Sharp 198 Methods Chomsky Resp. Intel. Pacifist Convergence: Anarchist ++ Accelerationist ++ Pacifist = PAA
Figure 3: Chronology of the three intellectual traditions that converge in peaceful anarcho-accelerationism (PAA): the anarchist tradition (from Godwin through Rojava), the accelerationist tradition (from Bookchin through Srnicek and Williams), and the pacifist tradition (from Tolstoy through Gandhi, Sharp, and Chomsky).
Anarcho-accelerationism rests on four intellectual pillars that bridge the historical and the contemporary. Kropotkin’s Mutual Aid (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006). Ostrom’s institutional analysis (Ostrom, 1990) demolished Hardin’s tragedy of the commons by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010). Following Stiegler (Stiegler, 2010), I adopt the concept of the pharmakon: every technology is simultaneously remedy and poison. This is not a reason for technological pessimism; it is a reason for technological governance (Feenberg, 2002). Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), I propose that the systems we build should embody the values they aim to promote.
3 Bridge-Building and Intersectional Liberation
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜5 summarizes the convergence.
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where I diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids (Farrell, 2016), self-hosted federated services that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives that eliminate dependence on extractive intermediaries (Scholz, 2016).
Socialism’s greatest insight is that material inequality undermines both freedom and democracy. Commons-based ownership, where resources are governed by their users according to Ostrom’s principles, distributes power more effectively than either state or market. The Mondragon cooperative corporation, with approximately 70,000 worker-owners and over € 11 billion in annual revenue, demonstrates that democratic enterprise is economically viable at industrial scale (Whyte and Whyte, 1991). The Emilia-Romagna region of Italy, where cooperative enterprises account for 30% of GDP, consistently ranks among the highest in quality of life in Europe (Restakis, 2010).
Environmentalism’s greatest insight is that economic activity must respect biophysical limits. The ecological crisis is structurally inseparable from the concentration of power: just 100 companies are responsible for 71% of global industrial greenhouse gas emissions (Griffin, 2017). Fossil fuels are inherently centralizing, while renewable energy is inherently decentralizing (Rifkin, 2011). Figure˜4 illustrates the dramatic decline in solar energy costs.
2,0102{,}0102,0122{,}0122,0142{,}0142,0162{,}0162,0182{,}0182,0202{,}0202,0222{,}02200.10.10.20.20.30.30.40.40.50.5YearLevelized Cost of Energy (USD/kWh)Solar PV (utility-scale)Fossil fuel range (gas)
Figure 4: Global weighted-average levelized cost of energy for utility-scale solar PV compared to natural gas, 2010–2022. Data from IRENA (IRENA, 2023). Solar costs have declined by approximately 89% in twelve years.
Feminism’s greatest insight, particularly in its materialist and care-ethics strands, is that the work of sustaining life is the foundation of all economic activity, yet it is systematically devalued (Federici, 2012; Tronto, 1993). Anarcho-accelerationism places care at the center of its value system not as a rhetorical gesture but as a design principle. Automation should liberate time for care, not replace care with machines.
The hacker ethic, articulated by Levy (Levy, 2010) and formalized by Stallman’s four freedoms (Stallman, 2002), holds that information should be free, access to computers unlimited, and decentralization promotes liberty. Anarcho-accelerationism is the political extension of this ethic beyond software. If code should be free, why not seeds? (Kloppenburg, 2014). If protocols should be open, why not energy grids? (Farrell, 2016). The Fediverse, built on the ActivityPub protocol and comprising Mastodon, PeerTube, PixelFed, Lemmy, Funkwhale, and BookWyrm, represents the most complete existing implementation of the hacker ethic applied to social communication. It demonstrates that a social communication platform can operate at scale, hosting over 10 million users across thousands of independently operated servers, without advertising, without algorithmic manipulation, and without centralized ownership (Zulli et al., 2020). Each server is a sovereign community with its own rules, moderation policies, and funding model, yet all communicate through a shared open standard. The Fediverse is anarchism in practice: federated, self-governing, commons-based, and voluntary.
The concept of the common good has deep roots in Catholic social teaching (Pope Francis, 2015), Islamic economics, Buddhist economics (Schumacher, 1973), and indigenous cosmovisions (Gudynas, 2011). Anarcho-accelerationism speaks this language without requiring religious commitment. Crucially, it respects the freedom of faith not merely as a political concession but on principled epistemological and ontological grounds. As argued by Lumbreras and Garrido-Merchán (2024), the ontological scope of theology categorically escapes the epistemological reach of artificial general intelligence. Through the frameworks of Saint Teresa of Avila’s Interior Castle and Saint Augustine’s Confessions, they demonstrate that human interiority is irreducible to computational processes. Full automation and AGI do not threaten authentic religious life; they liberate time for it.
Table 5: Bridge-building: shared values between anarcho-accelerationism and major traditions.
Tradition Core Value Shared Principle Contribution
Liberalism Individual freedom Autonomy, self-determination Material infrastructure for genuine autonomy
Socialism Material equality Fair distribution, solidarity Distributed ownership without state bureaucracy
Environmentalism Planetary limits Sustainability, stewardship Decentralized tech inherently aligned with ecology
Feminism Care and dignity Centrality of reproductive labor Care as design principle, not afterthought
Hacker ethic Information freedom Open source, decentralization Universalization of openness beyond software
Faith traditions Common good Shared dignity, mutual aid Commons governance; freedom of faith preserved
Intersectionality Liberation from all oppressions Elimination of structural inequality UDR removes material basis of all axes of oppression
Cooperativism Democratic enterprise Worker ownership, participation Platform cooperatives, digital tools for democracy
The key insight is that anarcho-accelerationism does not require anyone to abandon their deepest commitments. It asks only that they follow those commitments to their logical conclusion, and discovers that the conclusions converge.
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. I argue that UDR, implemented through the decentralized robotic commons, constitutes the most comprehensive intersectional intervention yet proposed, because it eliminates the material basis of all axes of oppression simultaneously. UDR eliminates class by definition: when every person has direct access to goods and services from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves. Racial oppression loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. The gendered division of labor loses its material foundation when all instrumental labor is automated. When access to resources is decoupled from productive capacity, the distinction between able and disabled bodies loses its economic significance. Economic dependence on discriminatory employers is eliminated for LGBTQ+ individuals. Both the young and the elderly gain full access to the resources of the commons. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. Table˜6 summarizes these interventions.
Table 6: Intersectional liberation through UDR: how the robotic commons eliminates the material basis of each axis of oppression.
Axis of oppression Material mechanism UDR intervention
Class Ownership of means of production Commons ownership eliminates class distinction
Race/ethnicity Discriminatory access to employment, housing, education Universal access independent of market participation
Gender Devaluation of care labor, wage gap All instrumental labor automated; care freely chosen
Disability Exclusion from productive economy Access decoupled from productive capacity
Sexuality Economic coercion via discriminatory employers Material independence from discriminatory institutions
Age Marginalization of non-productive ages Universal access regardless of labor capacity
Nationality Citizenship-based access to resources Material well-being not tied to state membership
The structural insight is this: intersectional oppression is maintained through material scarcity administered through markets. When markets are superseded by the robotic commons under UDR, the material infrastructure of oppression is dismantled at its foundation. Cultural prejudices do not disappear automatically, but prejudice loses its institutional teeth: the capacity to deny resources, opportunities, and dignity on the basis of identity.
Abraham Maslow’s hierarchy of needs (Maslow, 1943) captures a fundamental insight about human motivation: when basic survival needs are unmet, the higher capacities of creativity, intellectual exploration, and self-actualization remain largely inaccessible. The Liberation Stack and UDR are designed to satisfy the first four levels universally, physiological needs through the energy, manufacturing, and food layers; safety through UDR within ecological boundaries; belonging through communication and governance commons; esteem through knowledge commons, thereby liberating every human being to pursue self-actualization. Figure˜5 illustrates the mapping.
Physiological: food, water, shelter, energy Safety: health, stability, ecological security Belonging: community, connection, mutual aid Esteem: contribution, recognition, knowledge Self-actualization: creativity, spirituality, love Layers 0–2: Energy, Manufacturing, Food UDR + Ecological constraints Layers 3, 5: Communication, Governance Layer 4: Knowledge commons Real freedom: the goal
Figure 5: Maslow’s hierarchy of needs mapped onto the Liberation Stack and UDR. Self-actualization is what remains when all other needs are met: creativity, spirituality, love, and the free development of each person’s unique potential.
The transition from the current system to the care-centered society is progressive, not revolutionary, and it leverages the very mechanisms of capitalist development to transcend capitalism itself. Automation is already happening, driven by the profit motive. Every robot deployed, every AI system that replaces a function, is a step toward full automation. Anarcho-accelerationism does not oppose this process; it redirects it. The transition mechanism is a progressive continuum from Universal Basic Income to UDR, illustrated in Figure˜6. In the first stage, governments introduce and expand UBI, funded by taxation of automated production. In the second stage, communities simultaneously build the Liberation Stack: energy cooperatives, fablabs, agroecological food systems, mesh networks. As commons infrastructure grows, an increasing share of material needs is met outside the market. In the third stage, when the robotic commons can autonomously produce all material goods, money becomes unnecessary. UBI is replaced by UDR. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.
Current system Negative liberty Wage dependence Market access only Progressive UBI Partial protection Cash transfers Market still dominant Expanded UBI + Commons Growing commons Reduced market Synthetic liberty Full UDR Real freedom Post-monetary Commons-governed Automation beginsCommons growFull automationProgressive protection of displaced workers at every stage
Figure 6: The progressive transition from the current system to full UDR. At each stage, workers displaced by automation are protected by progressively expanding mechanisms.
This progressive path embodies what I term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act) illuminates the inadequacy of the neoliberal conception of freedom. Van Parijs (1995) argues that real freedom requires the material conditions for genuine self-determination. Synthetic liberty is the positive freedom that emerges when commons-based infrastructure provides those conditions. It is synthetic in two senses: it is constructed through deliberate institutional and technological design, and it synthesizes the insights of multiple traditions.
4 The Liberation Stack, Accelerationist Ecologism, and the Pareto Frontier
Anarcho-accelerationism proposes a concrete layered architecture for sociotechnical systems that I term the Liberation Stack, illustrated in Figure˜7. Each layer is composed of existing, mature, open-source technologies.
Layer 5: Governance (Decidim, Loomio, Pol.is, LiquidFeedback)Layer 4: Knowledge (arXiv, Wikipedia, open-source AI)Layer 3: Communication (Mastodon, Matrix, guifi.net, Signal)Layer 2: Food (Agroecology, FarmBot, Open Food Network)Layer 1: Manufacturing (Fablabs, 3D printing, WikiHouse)Layer 0: Energy (Solar cooperatives, OpenEMS, microgrids) Material
sovereignty
Informational
sovereignty
Figure 7: The Liberation Stack: a layered architecture for sociotechnical sovereignty. Lower layers (0–2) provide material sovereignty. Upper layers (3–5) provide informational sovereignty. Layers reinforce each other compositionally.
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar arrays and wind turbines, battery storage, smart microgrid management via open-source controllers such as OpenEMS, and peer-to-peer energy trading. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018). Full automation requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). I term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017). Table˜7 summarizes the frontier energy technologies.
Table 7: Frontier energy technologies for accelerationist ecologism: demonstrated innovations enabling abundant clean energy for full automation.
Technology Key milestone Projected impact Decentralization potential
Perovskite-silicon tandem solar 34.85% efficiency (Green et al., 2023) Solar below \$0.01/kWh Very high: modular, local
Enhanced geothermal Fervo Energy: 400 MW plant (Tester et al., 2006) Geothermal anywhere on Earth High: community-scale
Nuclear fusion NIF ignition Q=4.13; SPARC 2026 (Creely et al., 2020) Near-unlimited baseload by 2035–2040 Medium: regional plants
Iron-air batteries Form Energy: 100+ hours, <<\$20/kWh (Kwon et al., 2024) Solves multi-day intermittency Very high: iron and air
Green hydrogen H2Pro: 95%+ efficiency (Chatterjee et al., 2021) Replaces fossil fuels in industry High: modular electrolyzers
Long-duration storage DOE \$1.76B investment (U.S. Department of Energy, 2024) Grid-scale seasonal storage High: geological sites
[…]
Distributed manufacturing eliminates dependence on global supply chains. The global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is operational. WikiHouse provides open-source housing designs (Parvin, 2013). Food sovereignty combines traditional ecological knowledge with computational tools: agroecological farming, precision agriculture via IoT sensors and drones, and the Open Food Network connecting local producers and consumers. Cuba’s agroecological revolution demonstrated that small farmers can produce the majority of a nation’s food with minimal fossil fuel inputs (Altieri and Nicholls, 2012).
Digital sovereignty requires control over communication infrastructure at every level: the physical layer through community mesh networks following the guifi.net model with 40,000+ nodes (Baig et al., 2015); the protocol layer through ActivityPub for federation and Matrix for real-time communication; the application layer through Mastodon, PeerTube, PixelFed, and Lemmy; and the privacy layer through end-to-end encryption. The Fediverse hosts over 10 million users across thousands of independently operated servers, governed by their communities without centralized control (Zulli et al., 2020). Its governance model maps directly onto Ostrom’s design principles: clear boundaries (each server defines its rules), proportional equivalence (operators bear costs proportional to use), collective choice (community governance of moderation), monitoring (transparent moderation logs), graduated sanctions (muting, suspension, defederation), conflict resolution (inter-server negotiation), minimal recognition of rights (ActivityPub as W3C standard), and nested enterprises (individual servers within the broader network). The contrast with centralized platforms is instructive: when Elon Musk acquired Twitter in 2022, a single individual’s decisions could transform the governance of a platform used by hundreds of millions, an event that would be structurally impossible in a federated architecture. Knowledge commons ensure that information serves everyone: arXiv hosts 2.4+ million papers, Wikipedia’s 60 million articles in 300+ languages represent the largest knowledge commons in human history. Democratic self-governance is implementable as software: Pol.is for consensus visualization (Small et al., 2021), Decidim for participatory budgeting (Barandiaran et al., 2024), LiquidFeedback for delegable voting.
Table 8: The Liberation Stack: layers, functions, and exemplary technologies.
Layer Domain Function Key Technologies
5 Governance Collective decision-making Decidim, Loomio, Pol.is, LiquidFeedback
4 Knowledge Information commons arXiv, Wikipedia, open-source AI
3 Communication Digital sovereignty Mastodon, Matrix, guifi.net, Signal
2 Food Nutritional sovereignty Agroecology, FarmBot, Open Food Network
1 Manufacturing Production sovereignty Fablabs, 3D printing, WikiHouse
0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
The accelerationist ecological thesis is supported by rigorous econometric evidence. The Environmental Kuznets Curve hypothesis (Grossman and Krueger, 1995; Panayotou, 1993) proposes an inverted-U relationship between income and environmental degradation. The most comprehensive systematic review, screening 11,500+ papers, confirms that absolute decoupling of GDP from emissions has been achieved in several developed economies (Haberl et al., 2020). Le Quéré et al. (2019) found that renewable energy deployment is the primary driver of sustained emissions reductions in 18 developed economies. Way et al. (2022) demonstrate that a fast transition to renewables would save approximately \$12 trillion compared to maintaining fossil fuels. Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions can permanently redirect innovation toward clean technologies. Combined with dematerialization (McAfee, 2019), the evidence strongly supports the accelerationist thesis.
[…]
The apparent opposition between degrowth and accelerationism is a false dichotomy when viewed from the perspective of anarchist pacifism. Both movements share the same foundational commitment: opposition to capitalist domination, commitment to ecological sustainability, and pursuit of human flourishing through non-hierarchical social organization. The difference lies in technological strategy, not political values. This can be formalized as a multi-objective optimization problem. Define two objectives: $f_{1}$, the ecological risk during transition (probability of catastrophic environmental damage from betting on immature technologies), and $f_{2}$, the human suffering during transition (poverty, deprivation, and inequality persisting until post-scarcity technologies mature). Digital ecological degrowth occupies one region of the Pareto front: lower ecological risk (fewer technological gambles, reduced material throughput) but higher and more prolonged human suffering (because post-scarcity technologies arrive later, if at all, and material deprivation continues in the interim). Peaceful anarcho-accelerationism occupies the other region: higher ecological risk (betting on frontier technologies whose environmental impacts are not fully characterized) but lower human suffering (because technological abundance arrives sooner, relieving material deprivation more rapidly). No point on this front dominates the other; both positions are Pareto-optimal given different weightings of the two objectives. Figure˜8 illustrates this relationship.
DegrowthAccelerationismAnarchist pacifism(shared substrate)Tech. progressCurrent Pareto frontShifted frontEcological risk during transitionHuman suffering during transition
Figure 8: The degrowth-accelerationism tradeoff as a Pareto front. Both positions share anarchist pacifism as substrate and are Pareto-optimal given different weightings of ecological risk vs. human suffering. Technological progress (dashed curve) shifts the entire front toward the origin, reducing both risk and suffering. A mature movement supports both strategies simultaneously.
The critical insight is that the Pareto front is not fixed. Technological progress shifts it. Every advance in clean energy, every improvement in recycling efficiency, every breakthrough in sustainable manufacturing moves the entire front toward the origin: less risk and less suffering simultaneously. Accelerationism, by investing in these technologies, shrinks the Pareto front itself. This reframing transforms the debate: degrowthists and accelerationists are allies sharing anarchist pacifist foundations (Kallis, 2018; Hickel, 2020; Latouche, 2009), differing only in their position along a Pareto-optimal tradeoff. A mature movement would support both strategies simultaneously, hedging ecological bets through conservation and sufficiency (degrowth wisdom) while accelerating clean technology deployment (accelerationist commitment). What is fundamental and shared is the anarchist pacifism that underlies both positions: the rejection of domination, the commitment to commons governance, and the insistence on peaceful means. The technological strategy is secondary to the political ethics.
The Liberation Stack is not merely a list of technologies; it is a dependency graph with explicit preconditions for each layer. No layer can be sustainably deployed without the layers below it in the dependency ordering. This section makes the preconditions and gate conditions explicit, so that communities attempting to build the stack can evaluate their readiness for each layer and plan their implementation sequence rationally. Table˜9 summarizes the dependency structure.
Table 9: Liberation Stack preconditions and gate conditions. Each layer requires specific preconditions to be met before sustainable deployment. The dependency column indicates which lower layers must be operational.
Layer Domain Dependencies Gate condition for deployment
0 Energy None Community achieves >>50% renewable self-generation capacity
1 Manufacturing Energy Access to clean energy + open-source hardware designs for local fabrication
2 Food Energy, partial Manufacturing Energy for controlled environment agriculture + fabrication of farming tools and sensors
3 Communication Energy Electricity for servers, routers, and mesh nodes; can deploy in parallel with Manufacturing
4 Knowledge Communication Connectivity for accessing and sharing knowledge commons; functional federated platforms
5 Governance Communication, Knowledge Informed, connected citizens capable of sustained deliberation; digital literacy
Energy is the precondition for everything. Without clean energy sovereignty, no other layer achieves long-term sustainability: manufacturing depends on power for CNC machines and 3D printers, food production depends on power for irrigation pumps and controlled-environment systems, communication depends on electricity for servers and routers, and governance depends on the infrastructure that communication provides. The gate condition for the energy layer is that a community must achieve greater than 50% renewable self-generation capacity before the energy layer can be considered operational. This threshold is not arbitrary: below 50%, the community remains structurally dependent on external energy providers and vulnerable to supply disruptions or price manipulation. Current examples demonstrate that this gate is achievable. The World Bank is accelerating renewable deployment in developing countries through its Climate Change Action Plan, with particular emphasis on distributed solar in Sub-Saharan Africa and South Asia. The European Union’s Clean Energy Package includes directives enabling energy communities to generate, store, and sell renewable energy collectively, with over 3,400 energy cooperatives now operational across Europe (Kunze and Becker, 2015). Community solar programs in the United States have grown to over 6 GW of capacity, providing clean energy access to households that cannot install rooftop panels. The Just Energy Transition Partnerships (JETPs) negotiated by the World Bank and partner governments in South Africa, Indonesia, and Vietnam (World Bank, 2023) represent the most ambitious international framework yet for financing the transition from fossil fuels to renewables in developing economies, with combined pledges exceeding \$40 billion. A significant consideration within the energy layer is nuclear waste management. Conventional approaches to nuclear waste storage, including deep geological repositories, face both technical challenges and public opposition. An unconventional but increasingly plausible solution is space disposal via reusable rockets: as launch costs decline from approximately \$10,000/kg (Space Shuttle era) to projections below \$100/kg with fully reusable vehicles such as SpaceX Starship, the economics of launching vitrified high-level waste into solar orbit or toward the Sun become tractable. I note this not as a near-term proposal but as an illustration of how technological acceleration opens solution spaces that were previously dismissed as economically absurd.
Manufacturing requires the energy layer as a precondition. The gate condition is access to clean energy combined with open-source hardware designs sufficient for local fabrication of essential goods. Without energy, fablabs are expensive hobbies rather than productive infrastructure. With energy, the global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is operational and can produce everything from medical devices to housing components. Food sovereignty requires energy and partial manufacturing capability. The gate condition is energy for controlled-environment agriculture (vertical farms, greenhouses, aquaponics) combined with the capacity to fabricate farming tools, sensors, and irrigation systems locally. Communication requires only energy as a prerequisite and can be deployed in parallel with manufacturing. The gate condition is electricity sufficient to power servers, routers, and mesh nodes. The guifi.net model with 40,000+ nodes (Baig et al., 2015) demonstrates that community communication infrastructure can be built incrementally, with each new node extending the network. The Fediverse demonstrates that application-layer communication can be federated without centralized control (Zulli et al., 2020). Knowledge requires communication as a prerequisite: without connectivity, knowledge commons cannot be accessed or shared. The gate condition is functional federated communication platforms with sufficient bandwidth for knowledge exchange. Governance requires both communication and knowledge: informed deliberation presupposes access to information and the ability to communicate. The gate condition is a connected, informed citizenry with sufficient digital literacy to participate in online deliberative processes through tools such as Decidim (Barandiaran et al., 2024), Pol.is (Small et al., 2021), and Loomio.
The preconditions framework has a critical implication for the relationship between the Liberation Stack and the state. The framework proposes not revolution but progressive state dissolution: a gradual, layer-by-layer process in which the state’s functions are incrementally assumed by commons-based governance as each layer demonstrates competence. This process is explicitly incremental, reversible at every stage, and compatible with existing democratic institutions. It is not a seizure of state power; it is the progressive construction of alternatives that make state functions redundant.
In the first phase, the state remains fully operational but creates legal space for commons institutions. Energy cooperatives operate under existing cooperative law. Platform cooperatives register as conventional businesses. Community mesh networks operate under telecommunications regulations. The state’s role is to provide the legal framework within which commons institutions can form, experiment, and demonstrate viability. This phase requires no changes to constitutional order and is already underway in many jurisdictions: the EU energy communities directive, cooperative promotion legislation in most democracies, and digital commons policies in cities like Barcelona provide the legal infrastructure.
In the second phase, commons institutions provide an increasing share of public services, and the state’s role shifts from direct provider to guarantor of rights. Community energy cooperatives provide a growing fraction of the energy supply. Platform cooperatives handle increasing shares of local economic activity. Federated communication platforms provide alternatives to corporate social media. The state retains its regulatory and redistributive functions but delegates operational service delivery to commons institutions that have demonstrated competence. The key assumption is that commons institutions must demonstrate superior performance on measurable outcomes (cost, quality, accessibility, ecological impact) before the state delegates further functions to them.
In the third phase, federated commons governance handles the majority of resource allocation within participating communities, and the state’s role reduces to conflict resolution between federations and external defense. The governance layer of the Liberation Stack provides the institutional infrastructure for democratic self-governance at multiple scales: local communes for neighborhood decisions, regional federations for resource sharing, and inter-regional coordination for issues that transcend local boundaries. The state’s residual functions are those that require coercive authority: enforcement of ecological boundaries, protection of individual rights against community tyranny, and defense against external aggression.
In the fourth phase, the state withers as commons governance matures, and international federation handles coordination between self-governing regions. This phase is the most speculative and the most distant. It requires not only the successful operation of all Liberation Stack layers at regional scale but also the development of effective inter-regional governance protocols, the resolution of free-rider problems between federations, and the management of common-pool resources (atmosphere, oceans, biodiversity) at planetary scale. I present this phase as a theoretical horizon, not as a near-term prediction. The critical precondition for each phase is demonstrated competence at the previous level: if commons governance fails at local scale, there is no expansion to regional scale. If regional federations cannot coordinate effectively, there is no progression to inter-regional governance. This constraint is not merely prudential; it is structural. A commons governance system that has not proven itself at small scale has no basis for claiming authority at larger scale. The reversibility of each phase is essential: if a commons institution fails to deliver on its promises, the state’s functions can be restored without systemic collapse. This is the fundamental difference between progressive state dissolution and revolutionary state seizure: the former is incremental and reversible, the latter is sudden and irreversible.
The accelerationist ecological thesis is supported not only by the energy technologies already discussed but by a broader set of scientific frontiers that make the prospect of material abundance within ecological constraints increasingly credible. I survey these frontiers across four domains: econometric evidence for decoupling, biological engineering, advanced manufacturing toward post-scarcity, and complementary sciences.
The econometric evidence for the compatibility of technological progress with environmental sustainability has strengthened dramatically. The Environmental Kuznets Curve (EKC) hypothesis (Grossman and Krueger, 1995; Panayotou, 1993) proposes an inverted-U relationship between per capita income and environmental degradation: pollution rises during early industrialization but declines beyond a turning point as societies invest in cleaner technologies and environmental regulation. While Stern (2004) noted important statistical caveats regarding the generalizability of the EKC across pollutants and countries, the most comprehensive systematic review to date, screening 11,500+ papers and analyzing 835 full texts, confirms that relative decoupling of GDP from resource use and emissions is frequent, and absolute decoupling has been achieved in several developed economies for both production-based and consumption-based CO2 emissions (Haberl et al., 2020). The evidence has continued to strengthen. Le Quéré et al. (2019) analyzed 18 developed economies representing 28% of global emissions that achieved sustained emissions reductions between 2005 and 2015, finding that renewable energy deployment is the primary driver of decarbonization. The OECD Green Growth framework (OECD, 2011) documents that green productivity rose faster than traditional Total Factor Productivity across 30 member states. Way et al. (2022) demonstrate that technology cost forecasting based on Wright’s learning curves (Wright, 1936) and experience curves (Lafond et al., 2018) accurately predicts the cost decline trajectory of clean energy technologies, and that a fast transition to renewables would save the global economy approximately \$12 trillion compared to maintaining fossil fuel systems. Critically, Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions that redirect innovation toward clean technologies can permanently alter the trajectory of economic growth away from pollution-intensive paths. Combined with McAfee (2019)’s demonstration of dematerialization, the empirical finding that advanced economies are using fewer total resources even as GDP grows, and the World Bank’s framework for inclusive green growth (World Bank, 2012), the econometric evidence strongly supports the accelerationist thesis: the path to ecological sustainability runs through technological acceleration, not degrowth.
Biological engineering represents one of the most promising frontiers for simultaneously addressing energy, carbon capture, and food production. The RIPE Project (Realizing Increased Photosynthetic Efficiency) has achieved remarkable results: South et al. (2019) engineered a synthetic photorespiratory bypass in tobacco that increased field biomass productivity by over 40%, the largest photosynthetic improvement ever demonstrated in field trials. Kromdijk et al. (2016) accelerated recovery from photoprotection, increasing biomass by 14–20% through a mechanism common to all flowering plants, suggesting broad applicability across crop species. The C4 Rice Project at IRRI (Ermakova et al., 2020) aims to install C4 photosynthetic machinery into rice, potentially increasing efficiency by 50% in tropical conditions where rice is a staple for billions. Salesse-Smith et al. (2018) achieved a 30% increase in RuBisCO content and 15% increase in CO2 assimilation through enzyme overexpression in maize. Synthetic biology has created artificial photosynthetic systems that surpass natural ones: Miller et al. (2020) constructed an artificial chloroplast by encapsulating spinach photosynthetic membranes in cell-sized droplets powering the synthetic CETCH cycle, achieving CO2 fixation rates 100 times faster than previous approaches. The CETCH cycle itself (Schwander et al., 2016), comprising 17 enzymes from 9 organisms, uses an enzyme (CCR) that fixes CO2 approximately 20 times faster than plant RuBisCO. A particularly striking frontier is biophotovoltaics (BPV), which harness photosynthetic organisms to generate electricity directly. McCormick et al. (2015) review the field, documenting how cyanobacteria and algae transfer electrons extracellularly to electrode surfaces, producing electricity from sunlight and water with no fuel input. Plant-microbial fuel cells (Strik et al., 2008) generate electricity from living plant root exudates, with the Plant-e technology already commercialized for small-scale applications. Engineering CAM (Crassulacean Acid Metabolism) photosynthesis into C3 crops (Borland et al., 2014) could reduce water use by up to 80%, enabling food and bioenergy production on arid lands that are currently considered unsuitable for agriculture.
The frontier of advanced manufacturing is converging toward the ability to create virtually any material object from basic feedstocks with minimal waste. The relationship $E=mc^{2}$ implies that matter and energy are interconvertible; while wholesale matter creation from pure energy remains energetically impractical at current scales, pair production, the creation of matter-antimatter pairs from photon collisions, has been experimentally demonstrated. Breit and Wheeler (1934) provided the theoretical prediction, Burke et al. (1997) achieved the first laboratory demonstration at SLAC, and the STAR Collaboration at RHIC provided definitive evidence of the linear Breit-Wheeler process with real photons (STAR Collaboration, 2021). While pair production remains a physics demonstration rather than a manufacturing method, it establishes the fundamental principle that matter can be created from energy. In practical manufacturing, additive manufacturing is achieving near-zero-waste production: metal 3D printing achieves buy-to-fly ratios of 1:1 to 3:1, compared to 10:1 to 30:1 for conventional machining, dramatically reducing material waste (DebRoy et al., 2018). 3D bioprinting fabricates living tissues including skin, bone, vascular grafts, and cartilage (Murphy and Atala, 2014). DNA origami enables atomically precise construction at the nanoscale (Rothemund, 2006), progressing toward Drexler’s vision of molecular manufacturing (Drexler, 1992). Nuclear transmutation has created entirely new elements that do not exist in nature: using 48Ca beams on actinide targets, researchers at JINR, GSI, and RIKEN have synthesized elements 107–118 (Oganessian and Utyonkov, 2015), demonstrating that humans can convert one form of matter into another using kinetic energy at industrial research scales.
Several complementary scientific frontiers reinforce the accelerationist ecological thesis. In quantum computing for materials design, the variational quantum eigensolver (Peruzzo et al., 2014) and quantum chemistry algorithms (Cao et al., 2019) enable exponentially faster simulation of molecular behavior, potentially revolutionizing the rational design of catalysts, battery materials, and superconductors without costly trial-and-error synthesis. In nanotechnology, from carbon nanotubes (Iijima, 1991) to graphene (Novoselov et al., 2004), nanomaterials offer revolutionary capabilities for energy storage, water purification, and structural materials at a fraction of the weight and environmental cost of conventional alternatives. Nature’s 3.8 billion years of evolution provide what biomimicry researchers call open-source design blueprints for efficient buildings, drag-reducing surfaces, and ultra-tough materials (Benyus, 1997), democratizing innovation without intellectual property barriers. The circular economy demonstrates that waste can be eliminated by design: the cradle-to-cradle framework (McDonough and Braungart, 2002) and industrial symbiosis networks like Kalundborg (Jacobsen, 2006) show how inter-firm exchange of water, steam, and by-products can eliminate 635,000 tonnes of CO2 annually, consistent with anarchist principles of emergent, self-organizing order. In direct air capture, Keith et al. (2018) demonstrated industrial DAC at \$94–232 per tonne of CO2, and Hepburn et al. (2019) review ten pathways for utilizing captured CO2 in fuels, building materials, and chemicals, converting an atmospheric pollutant into feedstock for post-scarcity production. Finally, Drozdov et al. (2019) achieved superconductivity at 250 K ($23\degreeC$) in lanthanum hydride under high pressure, demonstrating that room-temperature superconductivity is physically achievable; practical superconducting power transmission could eliminate the 8–15% of electricity currently lost in grid distribution. Together, these scientific frontiers demonstrate that the technologies for post-scarcity without ecological destruction are not science fiction but active research programs with demonstrated results, declining costs, and clear pathways to deployment. The accelerationist ecological thesis is empirically grounded: we do not need to choose between human flourishing and planetary health. We need to accelerate the right technologies under commons governance.
Each layer of the Liberation Stack faces specific control structures that currently prevent commons deployment, and for each, concrete nonviolent strategies exist. At the energy layer, centralized utilities with regulatory capture resist community energy. The nonviolent response includes formation of energy cooperatives (Sharp’s method of alternative institutions), net metering advocacy (legal action), and community solar installations (constructive program). At the manufacturing layer, patent monopolies and proprietary supply chains resist open hardware. The response includes open-source hardware licensing, right-to-repair advocacy, and fablab networks as parallel infrastructure. At the food layer, agribusiness monopolies resist food sovereignty. The response includes community-supported agriculture, seed-saving networks, and agroecological education. At the communication layer, platform monopolies and surveillance infrastructure resist digital sovereignty. The response includes migration to the Fediverse (social noncooperation), community mesh networks (parallel infrastructure), and digital literacy campaigns. At the knowledge layer, academic publisher paywalls and proprietary AI resist knowledge commons. The response includes open-access mandates, community-trained AI models, and institutional repositories. At the governance layer, representative democracy captured by capital resists participatory governance. The response includes participatory budgeting campaigns, citizens’ assemblies, and prefigurative governance in cooperatives.
5 Social Mobilization: Nonviolent Strategy for the Commons Transition
The Liberation Stack provides the technical architecture for a care-centered society, but architecture without social force remains a blueprint. This section addresses the question of mobilization: what strategy would build the social force necessary to construct the Liberation Stack, and how can this be achieved through exclusively peaceful means?
The empirical evidence is unambiguous. Chenoweth and Stephan (2011), in the most comprehensive study of civil resistance ever conducted, analyzed 323 violent and nonviolent campaigns between 1900 and 2006 and found that nonviolent campaigns succeeded 53% of the time, compared to 26% for violent campaigns. Nonviolent campaigns succeeded even against the most brutal regimes, and the primary explanation is participation: nonviolent campaigns attract on average four times more participants than violent ones, because the barriers to participation (physical risk, moral objection, legal consequence) are dramatically lower. This finding has direct implications for the anarcho-accelerationist strategy: a movement that relies on technological commons-building, education, mutual aid, and democratic persuasion will attract far more participants than one that relies on confrontation.
Gene Sharp (Sharp, 1973, 2005, 2010) identified 198 methods of nonviolent action, organized into three categories: methods of protest and persuasion (public demonstrations, teach-ins, open letters, culture-jamming), methods of noncooperation (strikes, boycotts, data strikes, platform migration, tax resistance), and methods of nonviolent intervention (establishing alternative institutions, constructive programs, parallel governance, community infrastructure). The Liberation Stack maps naturally onto Sharp’s third category: each layer is a form of nonviolent intervention that creates alternative infrastructure rather than merely opposing existing power.
The question of consciousness, of how people come to understand their situation and choose to act, has been central to anarchist strategy since Godwin argued that rational education is the primary engine of social change. The anarchist tradition offers a rich repertoire: the Escuela Moderna of Ferrer i Guardia (Ferrer i Guàrdia, 1913) pioneered rationalist, anti-authoritarian education in early 20th century Spain. The workers’ athenaeums of the CNT combined cultural programming, literacy classes, and political education, reaching hundreds of thousands of workers before the Civil War. The Mujeres Libres organized 20,000 women through a combination of education, direct action, and mutual support (Ackelsberg, 2005). Updated for the digital age, consciousness-raising takes new forms. The precariat’s consciousness can be raised through exposure to the double nature of AI: its extraordinary potential for human liberation and its current deployment as what has come to be called AI slop, algorithmically generated low-quality content for advertising revenue, algorithmic management of gig workers, and surveillance-based behavioral manipulation. The pedagogical strategy is to show people what AI could do for them, automate drudgery, provide healthcare, generate knowledge, and then show them what AI actually does to them, target advertisements, manipulate attention, replace jobs without providing alternatives, generate deepfakes. AI coding agents could be community tools for building local software; instead they are deployed to replace programmers while enriching shareholders. AI medical diagnosis could provide universal healthcare; instead it is monetized behind paywalls. Agricultural AI could enable food sovereignty; instead it is used by agribusiness to optimize monocultures. This gap between potential and reality is itself the most powerful tool for consciousness-raising.
Syndicalism provides the organizational form for bottom-up pressure. The historical anarcho-syndicalist unions, the CNT in Spain, the CGT in France, the IWW in the United States, demonstrated that mass organizations built on direct democracy, mutual aid, and direct action could mobilize millions. The CNT’s achievement of 1.5 million members by 1936 was built not on ideological recruitment but on practical solidarity: mutual aid societies, workers’ schools, cooperative cafes, cultural centers. The lesson for the present is that the path to mass mobilization runs through practical benefit: communities adopt the Liberation Stack not because they are persuaded by anarchist theory but because community solar reduces their electricity bills, mesh networks provide free internet, fablabs let them repair their appliances, and governance platforms give them a voice in decisions that affect their lives. Ideology follows infrastructure, not the other way around.
Historical experience illuminates both possibilities and pitfalls. The CNT’s organizational genius reached 1.5 million members through practical solidarity, but its fatal error was joining the Republican government in 1936, contradicting its own anti-statist principles and subordinating the social revolution to the war effort (Bolloten, 1991). The Zapatistas demonstrated the successful use of communication technology for international solidarity while maintaining local autonomy: the internet transformed a regional indigenous movement into a global symbol of resistance (Castells, 2015). The indignados movement of 2011 in Spain demonstrated that mass horizontal mobilization could reshape political culture, and its institutionalization through Decidim represents one of the most successful translations of movement energy into participatory infrastructure (Barandiaran et al., 2024). The Occupy movement demonstrated strengths in consciousness-raising and prefigurative politics but weaknesses in constructive program: it articulated the problem of the 1% but did not build the alternative infrastructure to address it (Graeber, 2013).
The synthesis of these lessons yields a mobilization strategy with three components. First, education and fascination: show people the extraordinary potential of AI, robotics, and clean energy for human flourishing, and simultaneously expose the gap between this potential and its current deployment. This is not propaganda; it is the responsibility of intellectuals (Chomsky, 1967). Second, practical benefit: build Liberation Stack infrastructure that provides immediate material improvements to communities, from community solar to mesh networks to cooperative platforms. This is Corrêa’s social insertion (Corrêa, 2015): embedding anarchist principles within existing social needs rather than standing apart in ideological purity. Third, noncooperation and alternative institutions: as commons infrastructure grows, communities progressively withdraw from corporate dependencies, migrating from Twitter to Mastodon, from centralized utilities to community solar, from agribusiness to local food systems. Each withdrawal is a nonviolent act that simultaneously weakens corporate power and strengthens commons alternatives.
This strategy is not hypothetical. Contemporary initiatives across the globe demonstrate that commons-based mobilization is already operational at significant scale. The Just Energy Transition Partnerships (JETPs) negotiated in South Africa, Indonesia, and Vietnam (World Bank, 2023) represent the most ambitious international framework for financing the commons-based energy transition in developing economies, with combined pledges exceeding \$40 billion and explicit commitments to community ownership of renewable infrastructure. The International Labour Organization’s promotion of cooperative enterprises as vehicles for decent work and sustainable development provides an institutional framework at the global governance level. Barcelona’s digital sovereignty program has deployed Decidim as the city’s official participatory platform, enabling hundreds of thousands of citizens to participate in urban planning, budgeting, and policy deliberation (Barandiaran et al., 2024), demonstrating that Liberation Stack Layer 5 can operate within existing democratic institutions.
In the Global South, Kerala’s Kudumbashree network (Government of Kerala, 2023) is one of the largest women’s self-help and community development initiatives in the world, with approximately 4.5 million women organized into neighborhood groups, area development societies, and community development societies across all 1,070 panchayats in the state. Kudumbashree combines microfinance, collective enterprise, political participation, and mutual aid in a federated structure that maps remarkably onto the governance layer of the Liberation Stack. The Preston Model in the United Kingdom (Centre for Local Economic Strategies, 2019) demonstrates community wealth building through anchor institution procurement strategies: by redirecting the purchasing power of the city council, university, housing associations, and other anchor institutions toward local cooperatives and community enterprises, Preston has achieved measurable improvements in local employment, business formation, and economic resilience. In Jackson, Mississippi, Cooperation Jackson (Akuno and Nangwaya, 2017) is building a solidarity economy in one of the poorest cities in the United States, combining worker cooperatives, community land trusts, and participatory governance in a predominantly Black community, demonstrating that commons-based strategies can address racial and economic injustice simultaneously. The Evergreen Cooperatives in Cleveland, Ohio (Evergreen Cooperatives, 2023) apply a similar anchor institution strategy: worker-owned cooperatives in laundry, solar energy, and urban agriculture serve the procurement needs of major hospitals and universities in the University Circle neighborhood, creating living-wage jobs in communities with 40% poverty rates. These are not utopian experiments. They are operational programs with measurable outcomes, replicable models, and growing institutional support.
6 Empirical Evidence: Commons That Scale
A framework built on care and cooperation may sound idealistic. The evidence, however, is extensive. I present six case studies spanning different domains, scales, and geographies, all demonstrating that commons-based, decentralized, cooperative systems already work at scale.
The Linux kernel, with over 20,000 contributors across 1,700+ companies and independent developers, powers 96% of the world’s top web servers, 85% of smartphones, and 100% of supercomputers (Linux Foundation, 2023). It is governed without a state, owned by no corporation, and maintained through voluntary collaboration under copyleft licensing.
With 60+ million articles in 300+ languages, written and maintained by 300,000+ active volunteer editors, Wikipedia is the most comprehensive encyclopedia in human history (Jemielniak, 2014). It operates on donations, employs minimal staff, and has developed sophisticated self-governance mechanisms that map closely onto Ostrom’s design principles.
Founded in 1956 in the Basque Country, Mondragon encompasses approximately 70,000 worker-owners across some 90 cooperatives, a university, and 15 research centers. With over € 11 billion in annual revenue, it is the seventh-largest business group in Spain (Whyte and Whyte, 1991). The pay ratio between highest and lowest earners is capped at 6:1, compared to 399:1 in large US corporations (Mishel and Kandra, 2021). Table˜10 presents comparative performance data.
Table 10: Performance comparison: cooperative enterprises vs. conventional firms.
Metric Cooperatives Conventional Source
Labor productivity +12% baseline Pencavel (2012)
Worker wages +9% baseline Pencavel (2012)
Employee turnover 30% lower baseline Pencavel (2012)
5-year survival rate 80% 41% ICA/CECOP
Pay ratio (top/bottom) 6:1 (Mondragon) 399:1 (US large) Mishel and Kandra (2021)
[…]
Since 2012, the predominantly Kurdish region of Rojava has implemented democratic confederalism inspired by Bookchin’s municipalism (Öcalan, 2011). Governance is organized through nested councils with mandatory gender parity (Knapp et al., 2016). A population of approximately 4 million governs itself without a state, in the middle of a war zone. If it works during a civil war, it can work in peacetime.
Founded in 2004 in rural Catalonia, guifi.net is a community-owned telecommunications network with 40,000+ operational nodes (Baig et al., 2015). Anyone can connect and extend the network, provided they respect the network commons license. It provides internet access to rural communities that commercial ISPs consider unprofitable, demonstrating that physical infrastructure can be organized as a commons.
The Fediverse, built on the ActivityPub protocol standardized by the W3C, represents the most significant experiment in federated social communication in the history of the internet. Created by Eugen Rochko in 2016, Mastodon grew explosively during the crises of centralized platforms, particularly following the acquisition of Twitter by a single individual in 2022, an event that demonstrated with devastating clarity the fragility of centralized communication architecture. As of 2025, the Fediverse hosts over 10 million registered users across thousands of independently operated servers, spanning Mastodon for microblogging, PeerTube for video, PixelFed for images, Lemmy for discussion forums, Funkwhale for music, and BookWyrm for reading communities (Zulli et al., 2020; Gehl and Zulli, 2024). Each server is a sovereign community: it sets its own rules, elects its own moderators, chooses its own funding model (typically donations or cooperative membership), and decides which other servers to federate with or defederate from. No central authority exists. No algorithm decides what users see. No advertising revenue model distorts the incentive structure. The Fediverse demonstrates that social communication at scale does not require surveillance capitalism, algorithmic manipulation, or centralized ownership. It is the Liberation Stack’s Layer 3 in operation: digital sovereignty through federated, commons-based infrastructure. Its limitations are real, discoverability is harder than on algorithmic platforms, moderation in a federated environment faces novel challenges, and network effects still favor centralized incumbents, but its existence as a functioning, growing system constitutes empirical evidence that the anarchist architecture works for social communication.
Table˜11 summarizes the cross-case evidence.
Table 11: Empirical cases: commons-based systems at scale.
Case Scale Domain Duration Key Lesson
Linux 20,000+ contributors Software 33 years Peer production outperforms corporations
Wikipedia 300,000+ editors Knowledge 25 years Volunteer commons sustain quality at scale
Mondragon $\sim$}70,000 worker-owners Industry 70 years Democratic enterprise is economically viable
Rojava 4 million people Governance 14 years Federated democracy works at population scale
guifi.net 40,000+ nodes Telecom 22 years Physical infrastructure can be commons-based
Fediverse 10+ million users Communication 8+ years Federated social media operates without ads or surveillance
A common objection to transformative proposals is: who loses? In zero-sum frameworks, every gain implies a loss. Anarcho-accelerationism argues that the current system is negative-sum: it destroys more value than it creates through ecological degradation, social fragmentation, health costs of inequality, and the waste of human potential in meaningless labor (Graeber, 2018). The transition to commons-based infrastructure is positive-sum: everyone gains, including those who currently benefit from the status quo. Workers gain meaningful work, reduced hours as automation handles drudgery, democratic participation in economic decisions, and dignified conditions; the evidence from cooperatives shows 12% higher productivity, 9% higher wages, and 30% lower turnover compared to conventional firms (Pencavel, 2012). Entrepreneurs gain dramatically reduced fixed costs through open-source infrastructure, community energy, federated networks, and open governance tools (Bauwens et al., 2019). Communities gain resilient local economies through local energy production, local food systems, and local manufacturing, with a local economic multiplier 2 to 3 times higher than conventional businesses (Mitchell, 2006). The planet gains through decentralized renewable energy, local food systems, circular manufacturing, and repair culture that reduce emissions, waste, and resource extraction by design rather than as a costly afterthought (Raworth, 2017). Scientists and engineers gain through open-access publishing, open-source tools, open data, and collaborative research that accelerate discovery: open-access papers receive 18% more citations (Piwowar et al., 2018) and open-source scientific software is more reproducible and trustworthy (Stodden et al., 2014). Children and future generations gain a stable climate, preserved biodiversity, reduced inequality, and strong communities, the greatest inheritance any generation can leave.
Several objections deserve direct response with evidence-based counterarguments. The first objection is that this is utopian. The case studies in this section are not utopian: they are operational systems with decades of track record. Linux is not naive; it runs the internet. Mondragon is not utopian; it employs 70,000 people. The Fediverse is not a fantasy; it hosts 10 million users. Calling a working system utopian is not an argument; it is a refusal to engage with evidence. As Ursula K. Le Guin wrote: we live in capitalism, and its power seems inescapable, but so did the divine right of kings (Le Guin, 2014). The second objection is that commons-based systems do not scale. Linux scales to billions of devices. Wikipedia scales to hundreds of languages. The internet itself, the largest cooperative infrastructure ever built, scales to 5 billion users. The question is not whether commons-based systems scale but whether we are willing to invest in scaling them with the same resources we invest in proprietary alternatives. The third objection is that people are inherently selfish. This is empirically false. Kropotkin (Kropotkin, 1902), Ostrom (Ostrom, 1990), Henrich (Henrich, 2016), and extensive experimental economics research (Fehr and Fischbacher, 2003) demonstrate that humans are conditional cooperators: they cooperate when institutions are designed for cooperation. The tragedy of the commons is not a tragedy of human nature; it is a tragedy of institutional design. The fourth objection is that only capitalism can innovate. The internet was developed by public universities and government research labs, not venture capitalists. Linux was developed by volunteers. Wikipedia runs on donations. The most foundational technologies of the digital age, including TCP/IP, HTTP, HTML, SMTP, and DNS, are open protocols developed through collaborative, non-market processes. As Mazzucato (2013) demonstrates, capitalism is skilled at enclosing technological commons, not at creating them. The fifth objection is that free riders will destroy the commons. Ostrom’s design principles specifically address the free-rider problem through monitoring, graduated sanctions, and community enforcement. In open-source communities, social reputation, the intrinsic motivation of meaningful work, and the practical benefits of participation provide sufficient incentives (Lakhani and von Hippel, 2003). The more fundamental point is that in a commons-based economy of abundance, free riding is far less of a problem than in a scarcity economy: if the marginal cost of an additional user is near zero, as with digital goods, exclusion is more expensive than inclusion.
The Liberation Stack implies a research agenda for computer science that is both technically challenging and socially significant. The ActivityPub protocol, Matrix, and the broader Fediverse demonstrate that federated architectures can replace centralized platforms. However, significant challenges remain in moderation at scale, discoverability, identity management, and resistance to bad actors. Open problems include scalable content moderation in federated networks where no central authority can enforce global policies, decentralized reputation systems that resist both Sybil attacks and capture by powerful nodes, and privacy-preserving identity federation that enables accountability without surveillance. The current trajectory of AI development concentrates power in a handful of corporations with sufficient capital for training frontier models. Open-source AI represents a critical countermovement, but open weights alone are not sufficient: we need community-governed AI, models trained on consented data, auditable by their users, deployable locally, and designed to augment human judgment rather than replace it (Crawford, 2021). Open problems include efficient fine-tuning for community-specific needs, federated learning that preserves privacy, interpretable AI for democratic accountability, and energy-efficient inference for edge deployment. Pol.is, Decidim, and Loomio demonstrate that computational tools can enhance democratic deliberation, but current tools are rudimentary compared to what is technically possible. Open problems include scalable deliberation algorithms, fairness-aware voting mechanisms, AI-assisted policy simulation, and resistance to manipulation in online democratic processes. Community mesh networks like guifi.net demonstrate an alternative to corporate ISPs, but face challenges in bandwidth, reliability, and ease of deployment. Open problems include self-organizing mesh protocols for urban environments, quality-of-service guarantees in community networks, hybrid architectures bridging mesh and backbone networks, and user-friendly deployment tools that enable non-technical community members to extend the network.
7 Roadmap, Limitations, and Conclusion
A framework without a path of implementation is merely an aspiration. Table˜12 summarizes a four-phase roadmap toward the care-centered society. Crucially, the roadmap does not specify absolute dates. The transition can only begin when a sufficient base of social support has been built through sustained education, public deliberation, and demonstrated pilot successes, so that communities choose to adopt commons-based infrastructure through informed, voluntary, democratic decision-making. This precondition is not incidental but fundamental: a peaceful transition requires that the population freely decides to undertake it, and this in turn requires a prior process of education and consciousness-raising that cannot be scheduled in advance. The durations indicated below are therefore relative estimates, contingent on the pace at which this social foundation is established.
Table 12: Phased roadmap toward the care-centered society. Durations are relative estimates from the onset of each phase; absolute timing depends on the prior achievement of sufficient social support through education and democratic deliberation.
Phase Duration Core milestone Key deliverables
I. Seeds $\sim$}5 years Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities; launch federated UDR pilots at municipal scale
II. Roots $\sim$}10 years Regional infrastructure Interconnect pilots into regional commons federations; achieve energy and food sovereignty in participating regions
III. Canopy $\sim$}15 years Full automation Robotic systems autonomously manage energy, manufacturing, agriculture, logistics; UDR operational at regional scale
IV. Forest Open-ended Care-centered society Post-labor society across federations; human activity centered on care, creativity, education, art, science, community
Phase I (Seeds, approximately 5 years) begins once sufficient social support has been established through education, pilot demonstrations, and democratic deliberation. The technologies described in the Liberation Stack are already mature, the governance models are proven, and the cost curves for renewable energy and robotic hardware continue to decline. The task of this phase is integration: assembling existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets. The nonviolent mobilization strategy in this phase emphasizes education, fascination with AI’s potential for human liberation, and practical benefit through pilot Liberation Stacks that provide immediate material improvements. The gate condition for progression to Phase II is the successful operation of complete Liberation Stacks in at least 100 pilot communities across at least 10 countries, with measurable improvements in energy costs, food security, digital connectivity, and democratic participation relative to comparable non-participating communities. If Phase I does not achieve these metrics within the pilot communities, Phase II cannot begin. This is not arbitrary conservatism; it is the empirical validation without which scaling would be reckless.
Phase II (Roots, approximately 10 years) scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales? The nonviolent mobilization strategy in this phase shifts to noncooperation with corporate dependencies and the growth of federated commons alternatives: communities progressively withdraw from centralized utilities, corporate platforms, and extractive supply chains as commons infrastructure provides viable substitutes. Progressive state dissolution in this phase means the state creates formal legal frameworks for commons federations, recognizes inter-community governance protocols, and begins delegating operational service delivery to commons institutions that have demonstrated competence. The gate condition for progression to Phase III is energy and food sovereignty in participating regions (defined as greater than 80% of energy and food needs met by commons infrastructure), functional federated governance across at least 10 regional federations, and measurable reductions in Gini coefficient and ecological footprint relative to non-participating regions.
Phase III (Canopy, approximately 15 years) is the inflection point. When robotic systems can autonomously manage the entire material production chain, from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance, the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems. The nonviolent mobilization strategy in this phase emphasizes institutional consolidation: developing the governance protocols for post-monetary commons economies, negotiating resource-sharing agreements between federations, and establishing the ecological monitoring systems that enforce planetary boundaries as first-order constraints on production. Progressive state dissolution in this phase means federated commons governance handles the majority of resource allocation, and the state’s role reduces to conflict resolution between federations and external defense. The gate condition for Phase IV is the sustained autonomous operation of the complete material production chain across multiple federations, with ecological footprint demonstrably within planetary boundaries, for a period of at least five years.
Phase IV (Forest, open-ended) is the horizon. It represents not a fixed destination but a sustained process of institutional and cultural development. In a society where instrumental labor is no longer required, the central design challenge becomes the creation of institutions that support meaningful human activity: care work, education, scientific research, artistic creation, community governance, and the cultivation of interpersonal relationships. The reallocation of human time from instrumental to intrinsically motivated activity constitutes the substantive content of the framework’s normative vision. Progressive state dissolution in this phase means the state withers as commons governance matures, and international federation handles coordination between self-governing regions.
The Pareto front framework connects the roadmap to the degrowth-accelerationism debate. The phasing itself hedges against ecological risk: Phase I deploys only mature, well-characterized technologies (solar, wind, batteries), while frontier technologies (fusion, advanced geothermal, synthetic biology) enter the stack only as they demonstrate safety and sustainability. This phasing reflects the degrowth emphasis on caution with uncharacterized technologies, while the overall trajectory reflects the accelerationist emphasis on rapid clean technology deployment. The combination constitutes a rational response to a multi-objective optimization problem rather than a compromise between the two positions.
This roadmap rests on several explicit assumptions that I state transparently so that the framework can be evaluated against them. First, I assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results and as empirically observed over the past two decades. If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, I assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, I assume that open-source AI remains competitive with proprietary models. The current ecosystem of open-weight models suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, I assume that communities can develop effective commons governance at scale. The empirical evidence supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem. Fifth, I assume no catastrophic pandemic disrupts global supply chains during the critical scaling phases. The COVID-19 pandemic demonstrated that global supply chain disruptions can set back technological deployment by years; a pandemic of similar or greater severity during Phase II or III could significantly delay the roadmap. Sixth, I assume no major war disrupts the cooperative international environment necessary for the commons transition. The current geopolitical environment, including the conflict in Ukraine and tensions in East Asia, illustrates that great-power conflict remains a serious risk that could redirect technological investment from civilian to military applications and fragment the open-source ecosystem along geopolitical lines. Seventh, I assume that the political window for commons-based policy remains open: that democratic institutions continue to function sufficiently to enable cooperative legislation, energy community directives, and commons-friendly regulation. Democratic backsliding or authoritarian capture of state institutions would close this window. Eighth, I assume no catastrophic AI alignment failure: that frontier AI systems do not develop goals misaligned with human values in ways that resist correction. While I consider this risk low in the near term, its consequences would be severe enough to invalidate the entire framework.
The framework also operates within hard constraints that cannot be relaxed by any amount of technological progress. The most fundamental is ecological: the care-centered society must operate within planetary boundaries (Raworth, 2017). UDR is not unlimited consumption; it is universal access to what is needed for a dignified life, constrained by the carrying capacity of Earth’s ecosystems. The moral policy computed by DRL (Garrido-Merchán and Lumbreras-Sancho, 2023) must encode ecological sustainability as a first-order constraint, not an afterthought. A second constraint is material: full automation requires physical resources (rare earth minerals, silicon, lithium, copper) that are finite. Circular manufacturing, urban mining, and materials science innovation are necessary to close resource loops, but the total material budget of civilization is bounded. A third constraint is energetic: even with zero-marginal-cost renewable generation, the total energy budget of civilization is finite, and the robotic commons must allocate it efficiently.
I acknowledge several limitations of the framework as presented with full transparency. The timeline is inherently speculative; I offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not fully address geopolitical resistance from existing power structures: states, corporations, and military establishments have strong incentives to prevent the decentralization of productive capacity, and the history of social movements, from the Paris Commune to the Spanish Revolution, suggests that this resistance can be lethal. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised, potentially requiring significant restructuring of the automation framework. The UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; I present it here as a design principle and research direction, not as a finished policy proposal. The intersectional liberation thesis, while structurally sound, requires empirical validation: measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. The preconditions framework, while an improvement over vague roadmaps, relies on threshold values (50% renewable self-generation, 80% energy sovereignty) that are necessarily approximate and may require adjustment based on pilot experience. Finally, the progressive state dissolution framework assumes a cooperative relationship between state institutions and commons movements, which may not obtain in all political contexts: authoritarian states, captured regulatory agencies, or deeply polarized polities may resist commons-building regardless of its demonstrated benefits.
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the intersectional liberation thesis requires empirical validation: while the structural argument is sound, measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. Third, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? Fourth, the progressive transition from UBI to UDR requires pilot implementations in real communities, measuring not only technical performance but social outcomes including well-being, equity, ecological impact, community cohesion, intersectional equity, and Maslow-level satisfaction. Fifth, the accelerationist ecological thesis must be validated against planetary boundaries: can frontier clean energy technologies, deployed at the scale required for full automation, operate within Earth’s biophysical limits? Sixth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Seventh, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.
In summary, this paper has introduced peaceful anarcho-accelerationism as a rigorously defined sociotechnical framework, grounded it in a 200-year intellectual tradition, proposed a concrete technical architecture with explicit preconditions and gate conditions, and provided empirical evidence that commons-based systems already operate at scale across multiple domains. The framework’s principal contribution is the integration of automation theory, commons governance, intersectional analysis, and nonviolent strategy into a coherent design philosophy with testable hypotheses and a phased implementation roadmap. Significant empirical and institutional work remains before the framework’s more ambitious claims can be validated, but the existence of large-scale commons-based systems such as Linux, Wikipedia, Mondragon, Rojava, guifi.net, and the Fediverse demonstrates that the underlying organizational principles are not merely theoretical but operational. The framework is offered as a contribution to the interdisciplinary research agenda on the governance of automation, and its hypotheses, assumptions, and limitations have been stated explicitly to facilitate critical evaluation and refinement by the scholarly community.
Supplementary Material: Glossary of Categories and Terms
Term Definition
Peaceful Anarcho-Accelerationism A sociotechnical framework defined by seven analytical categories (ideology, strategy, social force, power, domination, self-management, social classes) for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care, achieved through exclusively nonviolent means.
Liberation Stack A six-layer architecture for sociotechnical sovereignty: Energy, Manufacturing, Food, Communication, Knowledge, Governance. Each layer uses existing open-source technologies and reinforces the others compositionally.
Universal Desired Resources (UDR) A post-monetary design principle in which every person has direct access to goods and services from the robotic productive commons, subject only to ecological sustainability constraints. Supersedes Universal Basic Income.
Synthetic Liberty The positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. Synthetic in two senses: constructed (not natural) and synthesizing insights of multiple traditions.
Accelerationist Ecologism The conviction that ecological sustainability is best achieved through clean energy abundance and technological acceleration within planetary boundaries, rather than through degrowth.
Compositional Sovereignty The property of the Liberation Stack whereby each layer can be adopted independently, but layers reinforce each other, yielding greater sovereignty as more layers are deployed.
Care-Centered Society A society whose fundamental organizing principle is care for persons rather than production for profit, enabled by decentralized full automation.
Pharmakon Following Stiegler: every technology is simultaneously remedy and poison. The question is never whether to use technology but who controls it, for what purposes, and under what institutional arrangements.
Moral Policy Following Garrido-Merchán and Lumbreras: the optimal reinforcement learning policy that maximizes the common good subject to humanist constraints, proposed as a computational tool for helping communities reason about ethical trade-offs.
Prefigurative Politics The principle that systems should embody the values they aim to promote: tools for liberation must not reproduce domination, and platforms for democracy must be governed democratically.
Principle of Totality and Interdependence Following Corrêa: the methodological principle that analytical categories of anarchism (ideology, strategy, social force, etc.) are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole.
Theory/Practice Unity The requirement that intellectual analysis of domination be inseparable from the construction of alternatives: theory without practice is speculation, practice without theory is blind activism.
Thought/Action The insistence that the political level of theoretical analysis must be embedded within the social level of popular movements and concrete organizing.
Form/Content The principle that organizational form (federated, horizontal, commons-based) must reflect content (liberation, care, non-domination).
Social Vectors The specific social forces that carry the potential for transformation: the precariat, open-source developers, cooperative movements, agroecological movements, community energy movements, Fediverse communities.
Social Insertion Following Corrêa’s especifismo: the practice of embedding anarchist principles within existing social movements and their concrete needs, rather than standing apart in ideological purity.
Especifismo A strategy originating in South American anarchism, systematized by Corrêa, emphasizing the need for specific anarchist political organization embedded within broader social movements.
Universal Approximation Theorem The theorem (Cybenko, Hornik) that a neural network with one hidden layer can approximate any continuous function to arbitrary precision, establishing that no mathematical barrier prevents automation of any formalizable task.
Deep Reinforcement Learning A framework where agents learn optimal policies through interaction with environments, establishing that physical and cognitive tasks are in principle learnable by artificial agents.
Federated Architecture A system design where independent, self-governing nodes communicate through shared protocols without central control, as exemplified by the Fediverse and email.
ActivityPub The W3C-standardized protocol for federated social communication, underlying Mastodon and the broader Fediverse.
Fediverse The network of federated social media platforms (Mastodon, PeerTube, PixelFed, Lemmy, Funkwhale, BookWyrm) communicating via ActivityPub.
Commons-Based Peer Production Large-scale collaboration without property rights or hierarchical command, as exemplified by Linux and Wikipedia.
Platform Cooperativism The model of cooperative ownership applied to digital platforms, where users and workers govern the platform democratically.
Ostrom’s Design Principles Eight institutional conditions for long-enduring commons: clear boundaries, proportional equivalence, collective choice, monitoring, graduated sanctions, conflict resolution, minimal recognition of rights, nested enterprises.
Non-domination The principle that no person, institution, or algorithm shall exercise coercive power over another.
The Precariat Following Standing: the growing class of workers in precarious, insecure, platform-mediated employment, identified as a primary social vector for transformation.
Techno-feudalism Following Varoufakis: the contemporary system where platform monopolies extract rent from digital serfs through surveillance capitalism and algorithmic management.
Intersectionality Following Crenshaw and Collins: the analysis of how systems of oppression (race, gender, class, etc.) intersect to produce compound disadvantages irreducible to single dimensions.
Pareto Front In the degrowth-accelerationism context: the set of strategies where no improvement in ecological risk is possible without increasing human suffering, and vice versa. Both positions are Pareto-optimal.
Negative Liberty Following Berlin: freedom from interference. The neoliberal conception of freedom.
Positive Liberty Following Berlin: freedom to act, to realize one’s potential. Requires material conditions, not merely absence of coercion.
AI Slop Algorithmically generated low-quality content produced for advertising revenue, exemplifying the misuse of AI potential under surveillance capitalism.
Precondition / Gate Condition A specific, measurable criterion that must be satisfied before a Liberation Stack layer can be considered operational or before progression to the next roadmap phase can begin. Gate conditions ensure that scaling is grounded in demonstrated competence rather than aspirational projection.
Progressive State Dissolution The gradual, layer-by-layer process in which state functions are incrementally assumed by commons-based governance as each Liberation Stack layer demonstrates competence. Explicitly incremental, reversible at every stage, and compatible with existing democratic institutions.
Environmental Kuznets Curve The hypothesis, first established by Grossman and Krueger and named by Panayotou, that environmental degradation follows an inverted-U relationship with income: rising during early industrialization and declining as societies invest in cleaner technologies. Systematic reviews confirm absolute decoupling in several developed economies.
Pareto-Optimal Strategy A strategy from which no improvement in one objective is possible without worsening another. In the degrowth-accelerationism context, both positions are Pareto-optimal given different weightings of ecological risk vs. human suffering during transition.
Just Energy Transition A framework for financing and managing the transition from fossil fuels to renewable energy in ways that protect affected workers and communities. Operationalized through Just Energy Transition Partnerships (JETPs) in South Africa, Indonesia, Vietnam, and other developing economies.